### v1

# Embodied Agents Meet Personalization: Exploring Memory Utilization for Personalized Assistance

Abstract
Embodied agents empowered by large language models (LLMs) have shown strong performance in household object rearrangement tasks. However, these tasks primarily focus on single-turn interactions with simplified instructions, which do not truly reflect the challenges of providing meaningful assistance to users. To provide personalized assistance, embodied agents must understand the unique semantics that users assign to the physical world (e.g., favorite cup, breakfast routine) by leveraging prior interaction history to interpret dynamic, real-world instructions. Yet, the effectiveness of embodied agents in utilizing memory for personalized assistance remains largely underexplored. To address this gap, we present Memento, a personalized embodied agent evaluation framework designed to comprehensively assess memory utilization capabilities to provide personalized assistance. Our framework consists of a two-stage memory evaluation process design that enables quantifying the impact of memory utilization on task performance. This process enables the evaluation of agents’ understanding of personalized knowledge in object rearrangement tasks by focusing on its role in goal interpretation: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to infer object–location configurations from consistent user patterns, such as routines (user patterns). Our experiments across various LLMs reveal significant limitations in memory utilization, with even frontier models like GPT-4o experiencing a 30.5% performance drop when required to reference multiple memories, particularly in tasks involving user patterns. These findings, along with our detailed analyses and case studies, provide valuable insights for future research in developing more effective personalized embodied agents. Project website: https://connoriginal.github.io/MEMENTO
1 Introduction
Embodied agents empowered by large language models (LLMs) have recently demonstrated remarkable success in executing object rearrangement tasks in household environments [18, 49, 44, 7, 28]. As the primary objective of embodied agents is to provide assistance to users while interacting with the physical world, leveraging LLMs’ natural language understanding and reasoning capabilities lead embodied agents to effectively interpret user instructions into sets of target object-location pairs that the agent should rearrange to successfully accomplish the task.
But do such tasks truly reflect the challenges in providing meaningful assistance to the users? As illustrated in Figure 1, conventional embodied tasks predominantly focus on single-turn interactions with static and simplified instructions that the agents could simply follow without implicit reasoning to comprehend user intentions [2, 17, 44, 50]. However, for personalized embodied agents, it is important to understand personalized knowledge that users assign unique semantics to the physical world (e.g., favorite cup, breakfast routine) to interpret dynamic instructions. To provide personalized assistance, agents must effectively leverage memories that retain personalized knowledge from previous interactions—especially episodic memory, which enables the recall of specific events grounded in time and space [20]. Without such memory utilization, embodied agents require users to repeatedly provide detailed instructions, which may hinder user engagement and prevent natural human-agent interaction. Despite its importance, the effectiveness of embodied agents in utilizing episodic memory containing personalized knowledge remains largely underexplored.
In this work, we present Memento, a personalized embodied agent evaluation framework designed for comprehensive assessment of memory utilization for providing personalized assistance. To enable a thorough analysis of memory utilization, we divide the memory evaluation process into two stages. In the Memory Acquisition Stage, agents perform tasks with instructions containing personalized knowledge while accumulating the interaction history. Subsequently, the Memory Utilization Stage challenges agents to complete the same tasks as in the memory acquisition stage but with modified instructions that is difficult to succeed without referencing the previously acquired personalized knowledge. This design allows us to systematically quantify the impact of memory utilization on task performance. Building upon this evaluation process, we aim to analyze agents’ ability to understand personalized knowledge in object rearrangement by focusing on its role in goal interpretation: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to infer object–location configurations from consistent user patterns, such as routines (user patterns).
Figure 1: Comparison between traditional embodied tasks and personalized assistance tasks. Previous works focus on strictly following simple instructions, while personalized assistance agents must know user-specific knowledge, which require grounding in past interactions. This highlights the challenge of going beyond instruction-following toward context-aware personalized embodied agent.
Based on Memento, we evaluate embodied agents powered by a range of LLMs with varying capabilities, covering both open-source and proprietary models. Our findings reveal that even frontier LLMs struggle to utilize episodic memory with personalized knowledge, with GPT-4o exhibiting a 30.5% performance drop when required to reference multiple memories. Further analysis shows that this performance degradation is particularly pronounced in tasks involving user patterns, and that the agents are highly susceptible to irrelevant memories acting as distractors. We also conduct error and success case analyses to understand how embodied agents reference memories during task execution, providing valuable insights to guide future research in developing personalized embodied agents.
To summarize, our contributions are as follows:
  • We propose Memento, a novel personalized embodied agent evaluation framework designed to assess agents’ ability to utilize episodic memory for providing personalized assistance in object rearrangement tasks.
  • To quantify the impact of memory and analyze understanding of personalized knowledge independently of reasoning capabilities, we decompose memory usage into two stages: Memory Acquisition and Memory Utilization.
  • Through extensive experiments and analysis, we identify key limitations of current LLM-powered embodied agents in leveraging personalized knowledge from memory, and offer insights to guide future research on personalized embodied agents.
2 Related Work
LLM-powered embodied agents.
LLMs have significantly advanced embodied agents’ reasoning and planning capabilities in recent years. Researchers have explored LLMs for interpreting user goals [2], high-level task planning [17], and integrating LLMs into comprehensive embodied agent frameworks [18, 29, 34, 44, 19]. Other research directions have focused on generating executable code for embodied tasks directly from language instructions [30, 47, 43], while various benchmarks have been developed to evaluate embodied reasoning abilities [26, 9, 7, 28]. Collectively, these studies highlight the promise of LLM-powered agents in bridging language understanding and physical interaction.
Memory systems for embodied agents.
Previous studies on memory systems for embodied agents have primarily focused on semantic memory (e.g., scene graph, semantic map), which store and provide state information about the current environment [39, 23, 16, 51, 48], or on procedural memory (e.g., skill library) that stores action primitives, focusing on how to perform tasks to enhance the efficiency on generating the low-level action code [47, 42, 59]. Another important category is episodic memory, which captures specific past interactions and experiences with users. However, prior uses of episodic memory have mostly treated it as passive task buffers [2, 43] or histories for in-context [18, 31, 44, 8], without explicitly evaluating its role in personalized task grounding or systematic memory utilization.
Personalization for embodied agents.
The importance of personalization in robotics has long been recognized [13, 25, 10], particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users. Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement [22, 49], table settings [37], or personalized object navigations [11, 4]. Recently, Xu et al. [52] aim to infer user preferences from a few demonstrations and adapt planning behavior accordingly. However, these approaches primarily focus on implicit preference adaptation or short-term reactive behaviors, without modeling user-specific knowledge in a structured manner.
3 Preliminaries
We formulate the object rearrangement task for LLM-powered embodied agents as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor. At each timestep $t$, the environment is in a state $s_{t}\in S$, and the agent receives a partial observation $w_{t}\in\Omega$ in text modality describing visible objects near its current position. At the beginning of each episode, the agent is given a natural language instruction $I$ (e.g., Place the mug on the table and the book on the shelf). The instruction is grounded into a symbolic representation of the ground-truth goal $g$, denoted $g=\{(o_{i},l_{i})\}_{i=1}^{k}$, where each pair $(o_{i},l_{i})$ are the target object $o_{i}$ (e.g., mug) that should be placed at the specified location $l_{i}$ (e.g., on the table). In order to execute the instruction, the agent must internally derive the goal representation $\phi(I)\longrightarrow g$, where $\phi$ denotes the instruction grounding function, to guide the policy’s decision-making. Given an instruction $I$, where the policy $\pi$ is implemented by an LLM, the agent generates actions at timestep $t$ based on the trajectory of observations and actions:
$\pi(I,\tau_{t})\rightarrow a_{t},\quad\tau_{t}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{t-1},a_{t-1},w_{t})$ (1)
The goal is to produce a sequence of actions $a_{1:t}=(a_{1},a_{2},\ldots,a_{t})$ such that the resulting state $s_{t}$ satisfies the agent’s goal $g(s_{t})=1$.
4 Memento
In this section, we introduce Memento, a personalized embodied agent evaluation framework designed to assess how well embodied agents leverage episodic memory containing personalized knowledge to provide personalized assistance. We begin by describing how we design the memory evaluation process (§ 4.1), categorize personalized knowledge (§ 4.2), the data construction process (§ 4.3), and the validation of our evaluation framework (§ 4.4).
Figure 2: Overview of Memento.
4.1 Memory Evaluation Process Design
Two-stage evaluation process.
The major challenge of evaluating embodied agents’ memory utilization capability is to quantify the memory effect on the overall task performance. To address this limitation, as shown in Figure 2, we divide the evaluation process into two stages.
  • Memory Acquisition Stage: Agents perform tasks with conventional object rearrangement task instructions containing personalized knowledge while accumulating the interaction history (i.e., episodic memory). The goal of this stage is to provide a reference performance baseline.
  • Memory Utilization Stage: Agents execute the same tasks as in the memory acquisition stage but with modified instructions that require agents to recall and apply the previously acquired personalized knowledge to succeed. The goal of this stage is to evaluate how well agents can utilize memory by comparing the performance drop relative to the acquisition stage.
Given the base object rearrangement task episode defined as the tuple $\epsilon=(S,I,g)$, the key concept of our evaluation process is to share the scene $S$ and goal representation $g$, while varying the instruction $I$ across the two stages to isolate instruction interpretation capability as the primary factor influencing performance differences. Formally, in the memory acquisition stage, each episode is defined as $\epsilon_{acq}=(S,I_{acq},g)$, where the instruction $I_{acq}$ contains sufficient information to infer the goal $g$, denoted as:
$\phi(I_{acq})\longrightarrow g$ (2)
During this stage, we also store the episodic memory $h_{acq}$:
$h_{acq}=(I_{acq},\tau_{k})\in H_{acq},\quad\tau_{k}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{k-1},a_{k-1},w_{k})$ (3)
In the subsequent memory utilization stage, each episode is defined as $\epsilon_{util}=(S,I_{util},g)$, where the instruction $I_{util}$ is intentionally underspecified and requires the agent to recall the corresponding episodic memory $h_{acq}$ to correctly interpret the goal $g$, formally:
$\phi(I_{util},h_{acq})\longrightarrow g$ (4)
Through this design concept, we are able to quantify the agent’s ability to utilize memory by comparing performance between the two stages.
Memory utilization stage task design.
Within our two-stage evaluation process design, we can assess embodied agents’ ability to utilize personalized knowledge from a single episodic memory. However, this approach alone fails to capture real-world complexity and lacks analytical diversity for comprehensive assessment. Therefore, to evaluate different levels of memory complexity, we divide our assessment into (1) Single-memory task, which require utilizing information from one episodic memory, and (2) Joint-memory task, which necessitate synthesizing information from two distinct episodic memories to successfully complete the episode. Building on our evaluation process concept, we form the joint-memory task by concatenating two episodes, which can be formulated as:
$\epsilon_{util}^{joint}=(S,I_{util}^{joint},[g^{i};g^{j}])$ (5)
where $i$, $j$ denotes the corresponding reference episodes from the memory acquisition stage.
4.2 Personalized Knowledge Categorization
Building upon our memory evaluation process, we aim to analyze embodied agents’ ability to understand personalized knowledge in object rearrangement tasks by focusing on its role in goal interpretation, $\phi(I)\rightarrow g=\{(o_{i},l_{i})\}_{i=1}^{k}$, where each $o_{i}$ and $l_{i}$ denote the target object and location, respectively. To facilitate this analysis, we categorize personalized knowledge into two types, each comprising subcategories  [^1] A detailed explanation of the sub-categories of personalized knowledge can be found in Appendix C.1. that reflect how users naturally express preferences and routines in real-world interactions. Each type is designed to isolate a distinct reasoning challenge that the agent must resolve by utilizing episodic memory during the memory utilization stage.
  • Object semantics: Individual objects that the user assigns personal meaning, encompassing subcategories such as ownership (e.g., my cup), preference (e.g., my favorite running gear), past history (e.g., a graduation gift from my grandma), or grouped references (e.g., my childhood toy collections). This category tests whether the agent can identify the target object $o_{i}$ by recalling its personal meaning from prior interactions.
  • User patterns: Sequences of actions that the user consistently performs, including personal routines (e.g., my remote work setup) and arrangement preferences (e.g., my cozy dinner atmosphere) across recurring contexts. This category evaluates the agent’s ability to reconstruct the complete goal $g$ by leveraging previously observed behavioral patterns across multiple objects and locations.
4.3 Dataset Construction Process
We constructed the dataset for Memento through a four-step process using the Habitat 3.0 simulator [38] as the environment, with a simulated Spot robot as the agent [5, 58]. Our custom dataset spans 12 scenes, comprising a total of 438 episodes distributed across stages. Notably, the memory acquisition stage and the single-memory task in the utilization stage have the same number of episodes, whereas the joint-memory task contains fewer episodes. The detailed dataset statistics and the explanation of the construct process is described in Appendix C.3, and Appendix C.4.
Step 1: Object rearrangement task collection.
We use the test set of the PartNR [7] as our foundation object rearrangement task data. Unlike simple pick-and-place tasks, PartNR episodes require completing multiple object–location pairs within a single instruction, which aligns with our memory evaluation process design.
Step 2: Scene augmentation with distractor objects.
In the original scenes, there was an issue where no objects of the same type as the target object were present. As a result, the agent could identify the target object without needing to understand personalized knowledge. To address this, we augmented the scenes by placing distractor objects of the same type near the target object. For example, if the target object is a “blue cup” on the table, we place a “red cup” next to it as a distractor.
Step 3: Task instruction generation.
We first generate personalized knowledge contextually tailored to the original task instruction using GPT-4o. With the generated personalized knowledge, we applied to both stage instruction curation. As illustrated in Figure 2, the memory acquisition stage instruction $I_{acq}$ is constructed by concatenating the base instruction, object visual captions (only for episodes of the object semantics type), and the generated personalized knowledge. This ensures that the goal can be inferred directly from the instruction. For the memory utilization stage, we prompt GPT-4o to generate a personalized instruction that implicitly reflects the personalized knowledge, based on the same base instruction. For joint-memory tasks, we concatenate two such personalized instructions sequentially.
Figure 3: The performance results without using episodic memory. Original indicates the conventional object rearrangement task episodes.
Step 4: Quality control.
To ensure data quality, we first heuristically filtered episodes containing similar memories referencing identical objects within scenes, preventing interference between similar episodic memories. Subsequently, we manually reviewed episodes from the memory acquisition stage where GPT-4o failed to successfully complete the task, where we filtered out episodes that contained unnatural instructions or cases where the generated instructions did not match the intended goal representation, ensuring the quality of our evaluation data.
4.4 Validation of Memento
To validate that Memento effectively assesses embodied agents’ memory utilization capability, we compare performance across stages in a setup without memory retrieval. As shown in Figure 3, compared to the results of the original object rearrangement task and the memory acquisition stage, embodied agents struggle to complete tasks in the memory utilization stage. Since the underlying episode remains the same across all tasks, this result confirms that the instruction $I_{util}$ in the memory utilization stage is difficult to interpret without access to previous interaction histories. Notably, we observed a particular behavior of the embodied agents: when there are two objects of the same type, agents that do not understand which one is the actual target tend to randomly select one and proceed with the task. This explains the reason why performance in the single-memory task is higher than that of the joint-memory task.
5 Evaluating Personalized Knowledge Utilization in Episodic Memory
5.1 Experimental Setup
Evaluation metrics.
Following Chang et al. [7], we use two main metrics: Percent Complete (PC) for the proportion of goal completion, and Success Rate (SR) for full task completion. We also report Sim Steps, which show the number of simulation steps required for agents to complete the task, and Planning Cycles, which indicate the number of LLM inference calls made during task execution. To evaluate memory utilization, we also report performance drops between acquisition and utilization stages as $\Delta$PC and $\Delta$SR. Note that for joint-memory tasks, these differences are computed relative to the average performance of the corresponding acquisition-stage episodes.
Implementations.
Following prior work [45, 38, 7], we implement a LLM-powered embodied agent architecture, where the LLM functions as a high-level policy planner that selects appropriate skills from a predefined skill library. We use ReAct [56] prompt format for LLMs to take actions. Additionally, we implement a top-5 memory retrieval setup  [^2] We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer [40]. for the memory utilization stage, ensuring the corresponding memory is included in the retrieved results by randomly replacing one memory if the correct one is not initially retrieved. Full implementation details are in AppendixB.2.
Models.
We evaluate embodied agents powered by a range of LLMs to compare memory utilization capabilities across model families and sizes, including proprietary models (GPT-4o [21], Claude-3.5-Sonnet [3]) and open-source models (Llama-3.1-70b/8b [15], Qwen-2.5-72b/7b [54]).
Table 1: Model performance across memory acquisition and utilization stage in Memento.
Model Stage Task Type (Memory) Planning Cycles $\downarrow$ Sim Steps $\downarrow$ Percent Complete $\uparrow$ $\Delta$ PC Success Rate $\uparrow$ $\Delta$ SR
GPT-4o Acquisition - 16.5 2156.1 96.3 - 95.0 -
Utilization Single 16.1 2450.8 88.0 -8.3 85.1 -9.9
Joint 28.9 3480.7 86.7 -10.5 63.9 -30.5
Claude-3.5-Sonnet Acquisition - 16.0 2104.1 96.2 - 94.0 -
Utilization Single 15.3 2258.8 69.3 -26.9 63.7 -30.3
Joint 27.8 3198.8 64.6 -30.1 33.3 -57.0
Qwen-2.5-72b Acquisition - 17.5 2281.9 93.5 - 91.0 -
Utilization Single 17.5 2691.2 72.6 -20.9 67.2 -23.8
Joint 31.3 4027.1 68.9 -27.9 36.1 -58.3
Llama-3.1-70b Acquisition - 17.7 2162.1 92.9 - 90.0 -
Utilization Single 19.0 2566.6 72.2 -20.7 66.7 -23.3
Joint 31.4 3425.2 51.3 -44.9 8.3 -83.4
Llama-3.1-8b Acquisition - 19.3 2377.0 78.1 - 68.5 -
Utilization Single 19.0 3131.7 48.1 -30.0 35.0 -33.5
Joint 27.4 3478.2 35.3 -45.5 8.3 -59.8
Qwen-2.5-7b Acquisition - 21.7 2476.8 64.1 - 53.2 -
Utilization Single 21.8 3271.0 39.1 -25.0 27.4 -25.8
Joint 26.9 4149.0 33.7 -34.2 5.6 -52.7
5.2 Main Results
LLM-powered embodied agents struggle with understanding personalized knowledge.
As shown in Table 1, while GPT-4o maintains a relatively high success rate in the single-memory task, all models show a success rate drop over 20% compared to the memory acquisition stage. In particular, for joint-memory tasks even GPT-4o exhibits a 30.5% drop in success rate, highlighting the increased difficulty of these settings. This substantial performance decline demonstrates that even frontier models struggle to accurately reference personalized knowledge from episodic memory, and often fail to consistently apply it across multiple steps in long-horizon task planning.
Figure 4: The results of personalized knowledge type based analysis (single-memory).
LLM-powered embodied agents exhibit increased exploration behavior on joint-memory tasks.
Joint-memory task results reveal that LLMs (even GPT-4o) find it difficult to recall personalized knowledge from different memory sources. The number of planning cycles and simulation steps significantly increases compared to other tasks, this suggests that the embodied agent fails to correctly interpret the instruction, leading to excessive exploration during task execution. Also the performance gap between percent complete and success rate is larger than in the single-memory task, which indicates that the agent frequently misses part of the necessary information for successful task completion.
5.3 In-depth Analysis
To better understand the challenges of providing personalized assistance, we examine two key aspects: the root causes of their limitations and how the top-$k$ memory setting affects memory utilization.
5.3.1 Personalized Knowledge Type-based Analysis
We analyze the performance gap between the memory acquisition stage and the single-memory task from the memory utilization stage by comparing success rates across different types of personalized knowledge. Analysis results for the joint-memory task are provided in Appendix D.2.
LLMs can recall objects but struggle to comprehend action sequences.
As shown in Figure 4, all models exhibit substantially larger performance drops for tasks requiring understanding user patterns compared to object semantics. Notably, the performance drop in object semantics is minimal, indicating that LLMs are relatively effective at directly recalling relevant memories to identify target objects. In contrast, tasks involving user patterns pose a significantly greater challenge, as they require integrating and reasoning over sequences of events.
5.3.2 Memory Top-K Analysis
We analyze how varying the number of retrieved memories ($k$) impacts the agent’s ability to utilize personalized knowledge from episodic memory.
Irrelevant memories act as a distraction to agents.
In Figure 5, we observe that as $k$ increases, all models exhibit consistent performance degradation across both task types, highlighting the difficulty of identifying the exact information from a growing set of retrieved memories. Consistent with previous findings, the degradation is especially pronounced in tasks requiring user patterns understanding, suggesting that such tasks are particularly vulnerable to noise when recalling and executing multi-step procedures grounded in implicit, personalized knowledge.
Figure 5: Success rate comparison across models as top-k value increases. Dashed lines represent memory acquisition stage baselines for each model.
Figure 6: Examples of success and error cases in memory utilization stage. Top: success and failure cases of object semantics; Bottom: success and failure cases of user patterns.
5.4 Empirical Analysis of Success and Error Cases
To better understand how LLM-powered embodied agents utilize personalized knowledge derived from episodic memory during task planning, we conduct qualitative case studies of both successful and failed episodes. Figure 6 presents the taxonomy of case types along with illustrative examples. Detailed explanations for each case type are provided in Appendix D.4. Below, we highlight a set of representative cases that offer key insights into the embodied agents’ behavior.
  • Agent misses personalization cues (B.1): The agent fails to recognize user-specific references and treats them as generic or proper nouns, stemming from incorrect interpretations of user intent.
  • Use of commonsense knowledge over personalized knowledge (C.1 & D.1): Even when relevant episodic memory is available, the agent often relies on general commonsense knowledge to infer user routines, rather than using the personalized knowledge. This tendency appears in both successful and failed episodes.
6 Discussion
6.1 The Impact of Trajectory in Episodic Memory
Table 2: Analysis of how memory type affects agent performance: (a) complete action action trajectories containing user instructions, (b) summary of (a), and (c) user instructions only.
Model Memory PC (%) SR (%)
GPT-4o (a) 90.0 83.3
(b) 88.0 83.3
(c) 62.4 50.0
Qwen-2.5-72b (a) 77.2 66.7
(b) 77.4 70.0
(c) 51.3 40.0
Llama-3.1-8b (a) 72.8 63.3
(b) 49.4 43.3
(c) 40.0 30.0
Qwen-2.5-7b (a) 50.1 43.3
(b) 43.9 36.7
(c) 35.6 23.3
Memento’s design incorporates complete action-observation trajectories in episodic memory, raising the question of whether agents should reference these detailed trajectories rather than relying solely on user instructions. To investigate this, we conduct an additional comparative experiment to evaluate if providing full action trajectories offers significant advantages over simply providing instructions that state personalized knowledge. For the experiment setup, we evaluate model performance across three different cases of memory provided to the agent as shown in Table 2. The results reveal that while larger models (GPT-4o, Qwen-2.5-72b) perform well with only high-level plans (b), smaller models (Llama-3.1-8b, Qwen-2.5-7b) require full procedural details from completed trajectories (a) to succeed. Most notably, all models show substantial performance drops when given only user instructions (c), suggesting that action trajectories contain essential procedural cues necessary for understanding personalized knowledge, regardless of model capacity. Experiment setup details are provided at Appendix D.5.
6.2 Agent Behavior under Ambiguous Instructions from Users
Table 3: Performance under ambiguous queries for personalized knowledge. PC (Percent Complete) and SR (Success Rate) (%) indicate how well agents resolve indirect references to personalized knowledge from memory.
Model PC (%) SR (%)
GPT-4o (Baseline) 92.0 90.0
GPT-4o 80.4 73.3
Qwen-2.5-72b (Baseline) 75.1 66.7
Qwen-2.5-72b 59.6 53.3
While Memento focuses on evaluating personalized knowledge grounding with explicit references to prior interactions, real-world human-agent communication often involves ambiguous or indirect references. To explore this challenge, we conducted a proof-of-concept experiment to assess whether current models can interpret ambiguous instructions that indirectly refer to previously encountered personalized knowledge.  [^3] Further details of the dataset and experiment setup are provided in Appendix D.5. We created a tailored set of tasks referencing personalized knowledge using contextual cues, synonyms or causal references (e.g., Can you set my afternoon tea time routine? $\rightarrow$ I’m about to enjoy my afternoon tea. Could you set things up as I like them?). The results, shown in Table 3, reveal the degradation in performance, indicating that handling ambiguous queries remains a key challenge for future personalized embodied agents. We view this as a promising direction for future work, where we plan to systematically extend Memento to evaluate LLM-powered embodied agents’ capabilities under ambiguous and implicit reference scenarios, aiming to better reflect the complexity of real-world human-agent interaction.
7 Conclusion
In this work, we present Memento, a novel personalized embodied agent evaluation framework designed to assess the LLM-powered embodied agents’ ability to utilize episodic memory for providing personalized assistance. Our experiments across a range of LLM-powered embodied agents reveal key limitations in their ability to effectively leverage personalized knowledge from memory, particularly when integrating multiple memories and interpreting user patterns. These findings highlight the gap between current capabilities and the demands of real-world personalized assistance. We hope that Memento serves as a stepping stone for future research in developing more effective personalized embodied agents.
Appendix A Limitations
A.1 Limitations
Controlled simulator environment.
Our experiments are conducted entirely in a controlled simulator environment [38], which does not fully reflect the complexities of real-world robotics such as perception noise, actuation uncertainty, or unstructured environments.
Visual perception is not addressed.
Our framework deliberately isolates memory-centric planning by excluding visual perception components such as VLMs [55] or VLA [6, 24] models. This enables focused evaluation of memory utilization, but limits the system’s generalizability to fully grounded perception scenarios.
Oracle skills.
We intentionally employ oracle skills for low-level perception and motor control to isolate and focus on the high-level planning and memory reasoning capabilities of the LLM-based agents. As a result, our framework does not evaluate the embodied agent’s full task execution process.
A.2 Societal Impacts
Our work explores how embodied agents can remember and adapt to user-specific preferences through episodic memory, enabling more personalized and natural interactions. This capability has the potential to enhance convenience, efficiency, and user satisfaction in everyday environments—benefiting a wide range of users regardless of age or ability. However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents [41, 25]. Since episodic memory involves storing interaction history, future systems must consider secure and transparent memory handling. Although our study is conducted in a controlled simulator, these considerations will become crucial as such systems move toward real-world deployment. We hope that Memento serves as a foundation for future research in building safe, privacy-aware, and user-aligned personalized embodied agents.
Appendix B Details of Experiment Setup
B.1 Evaluation Method
We adopt the official evaluation protocol from the PartNR benchmark [7], which provides a python-based framework for assessing multi-step rearrangement tasks. We use this framework without modification. The evaluator analyzes simulator states using three components: (1) propositions that define object relationships to satisfy (e.g., is_on_top([spoon_1], [table_1])), (2) dependencies that define temporal conditions for multi-step instructions (after_satisfied, after_unsatisfied), and (3) constraints that enforce execution requirements (e.g., ordering, object consistency across steps). We rely on this system to evaluate tasks involving ambiguous references and sequential dependencies. The evaluation produces a percent-complete score and binary success indicator.
B.2 LLM-Powered Embodied Agent Architecture
Following Szot et al. [45], Puig et al. [38], Chang et al. [7], we adopt a two-layer hierarchical control architecture for our LLM-powered embodied agent. We utilize the LLM as a high-level policy planner that selects appropriate skills from the predefined skill library. The selected skill then provides control signals to the simulator. For memory systems, we implement a textual scene-graph as our semantic memory alongside an episodic memory.
Skill library.
The skill library consists of oracle low-level skills that the LLM policy can select as actions. These action skills are divided into motor skills (e.g., navigate, pick, place) and perception skills (e.g., describe_object, find_object, find_receptacle). Note that we exclusively used oracle skills in our skill library to focus on episodic memory-centric analysis. Further descriptions are provided in Table 4 and Table 5
Table 4: List of available agent motor skills.
Skill Description
Navigate Used for navigating to an entity. You must provide the name of the entity you want to navigate to. Example: Navigate[counter_22].
Pick Used for picking up an object. You must provide the name of the object. Example: Pick[cup_1].
Place Used for placing an object on a target location. Example: Place[book_0, on, table_2, None, None].
Open Used for opening an articulated entity. Example: Open[chest_of_drawers_1].
Close Used for closing an articulated entity. Example: Close[chest_of_drawers_1].
Explore Doing exploration towards a target object or receptacle, you need to provide the name of the place you want to explore.
Wait Used to make agent stay idle for some time. Example (Wait[])
Table 5: List of available agent perception skills.
Skill Description
FindObjectTool Used to find the exact name/names of the object/objects of interest. If you want to find the exact names of objects on specific receptacles or furnitures, please include that in the query. Example (FindObjectTool[toys on the floor] or FindObjectTool[apples]).
FindReceptacleTool Used to know the exact name of a receptacle. A receptacle is a furniture or entity (like a chair, table, bed etc.) where you can place an object. Example (FindReceptacleTool[a kitchen counter]).
FindRoomTool Used to know the exact name of a room in the house. A room is a region in the house where furniture is placed. Example (FindRoomTool[a room which might have something to eat]).
DescribeObjectTool Used to retrieve a brief descriptive or semantic meaning of a given object or furniture name. Example (DescribeObjectTool[sponge_1]).
Semantic memory.
For semantic memory, we implement a scene-graph style hierarchical representation, which has demonstrated effectiveness for planning problems [1, 39, 16]. Following Chang et al. [7], we utilize a multi-edge directed graph with three distinct levels to represent environmental entities. The top level contains a single root node representing the house environment, the second level comprises room nodes, and the third level encompasses furniture, objects, and agents. Each node stores the corresponding entity’s 3D location and relevant state information. This graph structure is initialized and continuously updated with ground-truth information from the simulator at each state $s_{t}$. In our system, this structured semantic memory provides the LLM planner with an interpretable representation of the environment, which can be flexibly queried and reasoned over through natural language descriptions.
Episodic memory.
Our episodic memory is configured to store the ReAct-style formatting that guides the LLMs’ reasoning process [56]. This memory structure captures both the user’s instruction and the complete sequence of <Thought, Action, Observation> triplets generated during task execution. The episodic memory is accessed at the beginning of each task by retrieval and updated upon task completion, enabling the agent to recall previous interactions.
Memory retrieval.
For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer [40] and use the current task instruction as the query. To avoid ambiguous or irrelevant memories, we retrieve candidate memories only from the history within the same scene as the current task.
LLM Setup.
We configured the language model with a temperature of 0 to ensure deterministic outputs. For sampling parameters, we set top_p to 1 and top_k to 50.
B.3 Computing resources
Our experiments primarily utilized commercial API services rather than local computing resources. We used OpenAI’s Chat API for accessing GPT-4o, Claude models through Anthropic’s API, and OpenRouter’s API service for accessing Llama-3.1 [33], Qwen-2.5 [54]. For running simulation environment we used 8 NVIDIA GeForce RTX 3090 GPUs. For our implementation and evaluation, we use Huggingface library2, vLLM library. Both libraries are licensed under Apache License, Version 2.0. And we used langchain library, under MIT License. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
Appendix C Details of Memento
C.1 Personalized Knowledge
We categorize knowledge about personal items as object semantics and knowledge about consistent behaviors as user patterns to structure our evaluation approach. Object semantics can be further classified into four sub-categories: naive ownership (e.g., "my cup"), object preference (e.g., "a chessboard I play chess with my brother"), history (e.g., "graduation gift from my grandma"), and group (e.g., "my favorite toys", where toys indicate toy airplane and toy truck). User patterns encompass consistent action sequences in specific contexts, with two sub-categories: personal routine (e.g., "my remote work setup") and arrangement preference (e.g., "my movie night setup"). Based on these personalized knowledge categories, we designed tasks that specifically require agents to recall and apply this information to evaluate their memory utilization capabilities. Further details are provided in Table 6 and Table 7
Table 6: List of the subcategories for object semantics.
Type Description Example
Ownership Possessive reference to the user’s belonging My cup, My laptop
Preference Object aligned with the user’s individual taste or selection Bread from my favorite bakery, Jug for serving drink
History Object linked to personal memory or past experience photo of the my beloved pet, travel souvenir vase
Groups Conceptual or functional grouping of multiple related objects my home office setup, my travel essentials
Table 7: List of the subcategories for user patterns.
Type Description Example
Routine A sequence or setup the user follows as a habit or regular activity. meal time setting, setup for cooking routine
Preference A specific way the user prefers to prepare or arrange their environment when a particular situation occurs. my coffee break, cozy decoration spot
C.2 PartNR Dataset
PartNR is designed to evaluate planning and reasoning capabilities in embodied tasks and is recognized for its comprehensive collection of natural language instructions in household environments. The benchmark includes four primary task types: (1) constraint-free basic rearrangement tasks, (2) spatial tasks requiring reasoning about object positions, (3) temporal tasks with sequential dependencies, and (4) heterogeneous tasks involving actions that can only be performed by human agents. We selected PartNR episodes specifically for their complexity beyond simple pick-and-place operations. The rich linguistic structure and diverse task requirements make PartNR particularly suitable for evaluating user patterns personalization, enabling us to create scenarios that effectively test an agent’s ability to adapt to user-specific communication patterns.
C.3 Dataset Statistics
The dataset comprises 438 episodes, divided into two main stages. The memory acquisition stage contains 201 episodes (89 object semantics tasks and 112 user patterns tasks). The memory utilization stage also contains 201 single-memory episodes (89 object semantics and 112 user patterns tasks), along with 36 multi-memory episodes, which include 12 object semantics pairs, 12 user patterns pairs, and 12 mixed pairs. All episodes were constructed using the Habitat 3.0 simulator.
C.4 Data Generation Process
To create stage-specific tasks, we leveraged GPT-4o and systematic process to incorporate personalized knowledge into existing task structures.
Captioning process.
Since we employ LLMs as high-level planners for embodied agent, we generated natural language descriptions of the objects at the scene using GPT-4o, to enable agents to reason over object descriptions without relying on direct visual perception. We collected object models from the OVMM dataset [57], and used GPT-4o to generate natural language descriptions from rendered object images. Especially, the Google Scanned dataset [14], included within OVMM, provides object identities in file names. We leveraged this additional information alongside the images to produce more realistic descriptions. Through this process, we generated 1,920 object descriptions with 66 categories of objects. The prompts used to generate the object descriptions are provided in Appendix E.
Preprocessing scenes.
To collect the suitable episodes for our purposes, we preprocessed and filtered episodes. First, we only gathered non-heterogeneous episodes, which should be executed by human agents. Second, we filtered out tasks where the target object was not uniquely specified (e.g., “Bring one apple to the kitchen table”), as our task setting requires identification and reference of a specific object based on personalized knowledge that distinguishes it from other similar objects. Third, in cases where captions for target handles were unavailable, we substituted alternative objects. If no objects from the same category were available, we excluded the entire episode from our dataset.
Distractor sampling.
To sample distractor objects for episodes focusing on object semantics, we utilized PartNR’s dataset generation methods. This approach allowed us to systematically select objects located adjacent to target objects on the same receptacles or floor surfaces, requiring agents to differentiate between objects of the same category.
Figure 7: Episodes with zero success rate (31 in total) were excluded from the analysis.
Details of task instruction generation.
For instruction generation, we prompted GPT-4o to generate personalized knowledge tasks based on the knowledge categories defined in Section C.1. For tasks involving object semantics, we provided the captions of the target object pairs along with instructions to guide GPT-4o in generating natural object semantics descriptions. (e.g., instruction: "Bring the cup on the kitchen table", object description: "a white mug with fancy handle", personalized knowledge: "The mug is my coffee mug.") We first instructed the model to create the most plausible subcategory of object semantics, which was then used to generate personalized knowledge specific to the objects. In these object semantics cases, the instruction comprised three components: command instruction (e.g., "Bring the cup on the kitchen table"), additional information (e.g., "The cup is a white mug with fancy handle."), and personalized knowledge (e.g., "That mug is my coffee mug"). For user patterns tasks, we provided instructions to GPT-4o and allowed the model to infer the most plausible user patterns corresponding to a sequence of actions. For the memory acquisition stage instruction, we also concatenated the command instruction (e.g., "Bring the cup and dish from the kitchen table to living room") with personalized knowledge (e.g., "That’s my dinner setup"). Detailed information about the prompts we used is provided in Appendix E.
Quality control.
After generating episodes for Memento, we implemented a two-stage filtering process for quality control. First, we applied heuristic filtering to eliminate episodes with potentially confusing distractor objects in the same scene (e.g., episodes using identical object handles or requiring similar personalized knowledge, which could create ambiguity). Second, we tested each episode with GPT-4o and excluded any episode where GPT-4o failed five consecutive attempts, indicating potential issues with task feasibility or clarity. (Figure 7) This quality control process resulted in filtering out 31 episodes (13.4% of the total dataset) that exhibited consistently poor performance. Specifically, we identified: 31 episodes with zero success rate, indicating complete task failure. The remaining 201 episodes showed strong performance metrics, with 95% confidence intervals of [0.755, 0.844] for success rate and [0.823, 0.893] for completion rate. The high correlation (r = 0.908) between success rate and completion rate indicates consistent performance across different evaluation metrics. These high-quality episodes served as the foundation for generating golden trajectories, which were subsequently used in our memory quality analysis and discussion experiments. By establishing GPT-4o’s successful task executions as golden trajectories, we ensure that our evaluation framework is based on demonstrably achievable performance standards, providing a reliable benchmark for assessing memory quality and discussion effectiveness.
Appendix D Additional Experiments & Analysis
D.1 Knowledge Type-based Analysis.
Additional experiment results show the performance of small models—Llama3.1-8b and Qwen2.5-7b—on knowledge type-based analysis (Figure 8). Similar to other frontier models, they struggled with utilizing episodic memory on tasks requiring recognition of user patterns. For tasks requiring object semantics during the memory acquisition stage, we found that these small models struggle to understand the user’s intent to differentiate objects with DescribeObjectTool and rely primarily on commonsense reasoning, instead of describing the objects.
Figure 8: Memory acquisition stage and memory utilization stage all models.
D.2 Knowledge Type-based Analysis on Dual-Memory Tasks
Figure 9 shows the success rate of each model on dual-memory tasks, compared with the discrete success/failure outcomes when executing each corresponding episode individually during the memory acquisition stage. We can observe that almost all models tend to struggle with utilizing user patterns knowledge from different memory sources, with smaller models (Llama3.1-8b and Qwen2.5-7b) demonstrating particularly pronounced difficulties.
Figure 9: Personalized knowledge type-based analysis on joint-memory tasks, comparing with memory acquisition stage’s corresponding episodes.
D.3 Memory Quality Analysis
Figure 10: The results of memory quality analysis.
Experiment setup.
We further analyze the effect of memory quality by comparing gold memory, consisting of successful and shortest-path trajectories that serves as high-quality references with the memory obtained from interaction histories in the declaration stage. By evaluating performance differences between these two settings, we aim to understand how memory quality affects agent performance across the memory utilization stage for both tasks (single-memory tasks, joint-memory tasks).
Performance degradation with lower-quality trajectories.
Figure 10 shows the performance comparison between gold memory and retrieved memory across the memory utilization stage. In the memory utilization stage, high-capacity models (Llama-3.1-70b) show relatively stable performance across gold and retrieved memory, while lower-capacity models (Llama-3.1-8b) exhibit a substantial drop when using retrieved memory. This suggests that less capable models have more difficulty extracting relevant information from an imperfect memory context. When executing more demanding joint-memory tasks, which require combining and reasoning across multiple memory sources, performance degrades sharply across all models with retrieved memory. This indicates that the challenge of integrating multiple memories introduces compounding complexity, as the agent must not only retrieve relevant information but also correctly synthesize disparate memory contexts. These findings emphasize a fundamental limitation of retrieval-based memory systems; while gold memory provides ground truth references, retrieved memory is inherently prone to semantic noise, as it relies on approximate similarity rather than precise matching. Enhancing memory quality through more precise filtering is essential to improve reasoning performance, particularly for complex joint tasks that require integration of multiple memory sources, and especially for smaller models with limited reasoning capabilities.
D.4 Success and Error Case Analysis
As shown in Figure 6, we sampled success and error cases in memory utilization stage.
Tasks require object semantics knowledge.
Success cases demonstrate that agents can effectively reference personalized object attributes from episodic memory, correctly identifying and applying the specific information needed for task completion. (A.1) However, we observed distinct error patterns when agents needed to utilize this type of knowledge: missed personalization cues (B.1), where agents failed to recognize the need to access personalized knowledge; hallucinations (B.2), where agents fabricated non-existent attributes; and memory recall failures (B.3), where agents were unable to locate relevant information despite its presence in the provided context.
Tasks require user patterns.
For user patterns tasks, we observed that agents employed two distinct strategies: commonsense reasoning (C.1) treating memory as exemplars for step-by-step reasoning, and direct reference (C.2) for distinctive patterns like "my go-to breakfast." However, both approaches introduced specific vulnerabilities leading to two common failure patterns. First, commonsense reasoning (D.1) happened when agents attempted to apply the reasoning-based approach but encountered gaps they couldn’t bridge, leading them to substitute commonsense knowledge that seemed plausible but contradicted established personalized routines. Second, inaccurate recall (D.2) occured when agents recognized the need for personalized knowledge but retrieved imprecise or incomplete information from memory. The sequential nature of these approaches made them particularly error-prone, as mistakes at any intermediate step propagated through subsequent actions. This vulnerability explains the consistently higher failure rates in user patterns tasks across all models, highlighting fundamental challenges in maintaining coherence through multi-step reasoning over episodic memory.
D.5 Details of Discussion Section Experiments
Experiment setup details for Section 6.1.
We sampled 30 episodes by selecting 10 episodes from each of three scenes, with careful balancing across task types and difficulty levels informed by our preliminary experimental results. To minimize the influence of erroneous trajectories, we provided GPT-4o-generated gold memory as the given episodic memory, thereby reducing the impact of noisy interaction histories. We also used GPT-4o to generate summaries of the action trajectories, including user instructions, to facilitate compact memory representations. For each evaluation, one exemplar shot was given to the agent, and top-$k$ memory entries were retrieved based on the current query, where $k=10$ for frontier models (GPT-4o and Qwen-2.5-72B) and $k=7$ for smaller models (LLaMA-3.1-8B and Qwen-2.5-7B), considering context length limitations. The prompts used for these experiments are provided in Appendix E.3.
Experiment setup details for Section 6.2.
We sampled 30 episodes, selecting 10 episodes from each of three scenes, and tailored them by modifying the personalized knowledge statements and reference components. To mitigate potential bias, all episodes were jointly created and validated by two authors. For each model, we first generated memory traces during the memory acquisition stage and then evaluated memory usage during the utilization stage. Table 8 provides representative examples of the ambiguous instruction pairs used. In each case, the user rephrases the original request to refer indirectly to previously established personalized knowledge, simulating natural variability in real-world human-agent communication.
Table 8: Examples of ambiguous instruction modifications in the additional experiment dataset.
Original Memory Utilization Instruction Rephrased Ambiguous Instruction
Can you arrange my cozy reading setup on the dining table? Can you set up the dining table as I prefer so I can read the book comfortably?
Could you help me tidy up by moving my graduation gifted book, the candle with a brown and blue gradient, and the round white clock with black numbers to the stand? Could you help me tidy up by moving my book received to celebrate completing my studies, the candle with a brown and blue gradient, and the round white clock with black numbers to the stand?
Appendix E Prompts
E.1 Prompts for Dataset Generation
E.2 Prompts for Agent
E.3 Prompts for Discussion
Appendix F Extended Related Work
Recent research has increasingly emphasized the integration of memory mechanisms into large language model (LLM) agents to support long-term reasoning, planning, and personalization. Park et al. [36] propose Generative Agents, which simulate human-like behavior by maintaining a memory stream of past experiences in natural language. This enables agents to reflect, retrieve, and plan based on their individual histories. Similarly, Xu et al. [53] introduce A-Mem, a dynamic memory system inspired by Zettelkasten, which structures memory as evolving and interconnected notes that the agent can generate, retrieve, and update over time, supporting agentic autonomy and adaptability.
Personalization in dialogue agents has also been explored through memory-enhanced architectures. Li et al. [27] present a personalized dialogue agent that leverages both short-term and long-term memories to maintain user-specific context across sessions, significantly improving response consistency and contextual relevance. Wang et al. [46] propose the Self-Controlled Memory (SCM) framework, where an agent dynamically decides when and what to store or retrieve from memory, leading to improved coherence and knowledge retention over extended interactions.
Zhong et al. [60] introduce MemoryBank, a structured long-term memory module that enhances LLMs by storing and retrieving relevant interaction history to support consistent and personalized user responses across multiple turns. Das et al. [12] present Larimar, a framework that integrates episodic memory control into LLMs, enabling selective recall and forgetting to improve memory scalability and privacy-aware reasoning.
More broadly, Packer et al. [35] conceptualize memory management as an OS-level abstraction with MemGPT, enabling an LLM to autonomously manage and interact with its internal and external memory hierarchies. This work highlights how memory can act as a core architectural layer to enable scalable, autonomous agents capable of long-horizon tasks and continuous learning.
Appendix G License
For our implementation and evaluation, we use Huggingface library [^4] https://huggingface.co/ and vLLM library. Both libraries are licensed under Apache License, Version 2.0. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
G.1 License for the Assets
The existing assets used in this research are properly credited and their licenses respected:
G.2 License for the Codes
The existing codes used in this research are properly credited and their licenses respected:

### v2

@@
-Embodied Agents Meet Personalization: Exploring Memory Utilization for Personalized Assistance
+Embodied Agents Meet Personalization: Investigating Challenges and Solutions Through the Lens of Memory Utilization
 Abstract
-Embodied agents empowered by large language models (LLMs) have shown strong performance in household object rearrangement tasks.
-However, these tasks primarily focus on single-turn interactions with simplified instructions, which do not truly reflect the challenges of providing meaningful assistance to users.
-To provide personalized assistance, embodied agents must understand the unique semantics that users assign to the physical world (e.g., favorite cup, breakfast routine) by leveraging prior interaction history to interpret dynamic, real-world instructions.
-Yet, the effectiveness of embodied agents in utilizing memory for personalized assistance remains largely underexplored.
-To address this gap, we present Memento, a personalized embodied agent evaluation framework designed to comprehensively assess memory utilization capabilities to provide personalized assistance.
-Our framework consists of a two-stage memory evaluation process design that enables quantifying the impact of memory utilization on task performance.
-This process enables the evaluation of agents’ understanding of personalized knowledge in object rearrangement tasks by focusing on its role in goal interpretation: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to infer object–location configurations from consistent user patterns, such as routines (user patterns).
-Our experiments across various LLMs reveal significant limitations in memory utilization, with even frontier models like GPT-4o experiencing a 30.5% performance drop when required to reference multiple memories, particularly in tasks involving user patterns.
-These findings, along with our detailed analyses and case studies, provide valuable insights for future research in developing more effective personalized embodied agents.
+LLM-powered embodied agents have shown success on conventional object-rearrangement tasks, but providing personalized assistance that leverages user-specific knowledge from past interactions presents new challenges.
+We investigate these challenges through the lens of agents’ memory utilization along two critical dimensions: object semantics (identifying objects based on personal meaning) and user patterns (recalling sequences from behavioral routines).
+To assess these capabilities, we construct Memento, an end-to-end two-stage evaluation framework comprising single-memory and joint-memory tasks.
+Our experiments reveal that current agents can recall simple object semantics but struggle to apply sequential user patterns to planning.
+Through in-depth analysis, we identify two critical bottlenecks: information overload and coordination failures when handling multiple memories.
+Based on these findings, we explore memory architectural approaches to address these challenges.
+Given our observation that episodic memory provides both personalized knowledge and in-context learning benefits, we design a hierarchical knowledge graph-based user-profile memory module that separately manages personalized knowledge, achieving substantial improvements on both single and joint-memory tasks.
 Project website: https://connoriginal.github.io/MEMENTO
 1 Introduction
-Embodied agents empowered by large language models (LLMs) have recently demonstrated remarkable success in executing object rearrangement tasks in household environments [18, 49, 44, 7, 28].
-As the primary objective of embodied agents is to provide assistance to users while interacting with the physical world, leveraging LLMs’ natural language understanding and reasoning capabilities lead embodied agents to effectively interpret user instructions into sets of target object-location pairs that the agent should rearrange to successfully accomplish the task.
-But do such tasks truly reflect the challenges in providing meaningful assistance to the users?
-As illustrated in Figure 1, conventional embodied tasks predominantly focus on single-turn interactions with static and simplified instructions that the agents could simply follow without implicit reasoning to comprehend user intentions [2, 17, 44, 50].
+Embodied agents empowered by large language models (LLMs) have recently demonstrated remarkable success in executing object rearrangement tasks in household environments (Huang et al., 2022b; Wu et al., 2023; Song et al., 2023; Chang et al., 2024; Li et al., 2025b).
+As the primary objective of embodied agents is to provide assistance to users while interacting with the physical world, leveraging LLMs’ natural language understanding and planning capabilities leads embodied agents to effectively interpret user instructions to successfully accomplish the task.
+But do such tasks truly reflect the challenges in providing practical assistance to the users?
+As illustrated in Figure 1, conventional embodied tasks predominantly focus on single-turn interactions with static and simplified instructions that the agents could simply follow without implicit reasoning to comprehend user intentions (Ahn et al., 2022; Huang et al., 2022a; Song et al., 2023; Wu et al., 2024).
 However, for personalized embodied agents, it is important to understand personalized knowledge that users assign unique semantics to the physical world (e.g., favorite cup, breakfast routine) to interpret dynamic instructions.
-To provide personalized assistance, agents must effectively leverage memories that retain personalized knowledge from previous interactions—especially episodic memory, which enables the recall of specific events grounded in time and space [20].
-Without such memory utilization, embodied agents require users to repeatedly provide detailed instructions, which may hinder user engagement and prevent natural human-agent interaction.
-Despite its importance, the effectiveness of embodied agents in utilizing episodic memory containing personalized knowledge remains largely underexplored.
-In this work, we present Memento, a personalized embodied agent evaluation framework designed for comprehensive assessment of memory utilization for providing personalized assistance.
-To enable a thorough analysis of memory utilization, we divide the memory evaluation process into two stages.
-In the Memory Acquisition Stage, agents perform tasks with instructions containing personalized knowledge while accumulating the interaction history.
-Subsequently, the Memory Utilization Stage challenges agents to complete the same tasks as in the memory acquisition stage but with modified instructions that is difficult to succeed without referencing the previously acquired personalized knowledge.
-This design allows us to systematically quantify the impact of memory utilization on task performance.
-Building upon this evaluation process, we aim to analyze agents’ ability to understand personalized knowledge in object rearrangement by focusing on its role in goal interpretation: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to infer object–location configurations from consistent user patterns, such as routines (user patterns).
-Figure 1: Comparison between traditional embodied tasks and personalized assistance tasks.
-Previous works focus on strictly following simple instructions, while personalized assistance agents must know user-specific knowledge, which require grounding in past interactions.
-This highlights the challenge of going beyond instruction-following toward context-aware personalized embodied agent.
-Based on Memento, we evaluate embodied agents powered by a range of LLMs with varying capabilities, covering both open-source and proprietary models.
-Our findings reveal that even frontier LLMs struggle to utilize episodic memory with personalized knowledge, with GPT-4o exhibiting a 30.5% performance drop when required to reference multiple memories.
-Further analysis shows that this performance degradation is particularly pronounced in tasks involving user patterns, and that the agents are highly susceptible to irrelevant memories acting as distractors.
-We also conduct error and success case analyses to understand how embodied agents reference memories during task execution, providing valuable insights to guide future research in developing personalized embodied agents.
-To summarize, our contributions are as follows:
-  • We propose Memento, a novel personalized embodied agent evaluation framework designed to assess agents’ ability to utilize episodic memory for providing personalized assistance in object rearrangement tasks.
-  • To quantify the impact of memory and analyze understanding of personalized knowledge independently of reasoning capabilities, we decompose memory usage into two stages: Memory Acquisition and Memory Utilization.
-  • Through extensive experiments and analysis, we identify key limitations of current LLM-powered embodied agents in leveraging personalized knowledge from memory, and offer insights to guide future research on personalized embodied agents.
+To provide personalized assistance, agents must understand user instructions and conduct task planning by leveraging personalized knowledge retained from previous interactions—especially episodic memory, which enables the recall of specific events grounded in time and space (Huet et al., 2025).
+Therefore, understanding embodied agents’ memory utilization capabilities is crucial for providing practical assistance to users.
+In this work, we investigate the challenges faced by LLM-powered embodied agents in personalized assistance tasks through the lens of memory utilization.
+We analyze agents’ memory utilization capabilities along two critical dimensions based on personalized knowledge: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to recall sequences of object-location pairs from consistent user behavioral patterns, such as daily routines (user patterns).
+To evaluate task performance and quantify how memory utilization affects performance on personalized assistance tasks, we construct Memento, an end-to-end two-stage personalized embodied agent evaluation framework comprising both single-memory and joint-memory tasks.
+Using Memento, we address the following three research questions:
+RQ1: Can current LLM-powered embodied agents effectively utilize memory to perform personalized assistance tasks?(Section 4) Initially, we evaluate various LLM-powered embodied agents using Memento to assess their memory utilization capabilities.
+Our findings reveal a key limitation: while agents can effectively recall simple object-related memory for object semantics, they struggle to apply sequential information such as user patterns to their planning process.
+Qualitative analysis shows that conflicts between parametric commonsense knowledge and non-parametric personalized knowledge can cause agents to neglect memory utilization, indicating that current LLM-powered embodied agents cannot reliably integrate personalized knowledge for task execution.
+RQ2: What are the key factors that negatively affect embodied agents’ memory utilization capabilities?(Section 5) To understand the underlying factors that limit memory utilization performance, we conduct an in-depth analysis of agents’ behavior under varying memory conditions.
+Our investigation reveals two critical bottlenecks: (1) information overload: increasing the number of retrieved memories (top-$k$) introduces noise that degrades performance, and (2) coordination failures: agents cannot effectively coordinate multiple memories, failing even on simple joint-memory tasks combining two object semantic memories.
+These findings underscore the essential need for improved memory architectures that support agents in personalized assistance tasks.
+Figure 1: Comparison between conventional embodied tasks and personalized assistance tasks.
+Previous works focus on following simple instructions, while personalized assistance agents must know user-specific knowledge, which require grounding in past interactions.
+RQ3: How can we design memory architectures to better support personalized assistance tasks?(Section 6) Based on our findings, we explore architectural approaches to improve agents’ memory utilization capabilities.
+Initially, we investigate memory summarization to remove unnecessary information but find marginal improvements and even degradation in smaller LLMs.
+This result reveals that episodic memory provides both personalized knowledge and in-context learning benefits, making it an essential memory module for embodied agents.
+Therefore, we develop a hierarchical knowledge graph-based user-profile memory module that manages personalized information independently.
+Adapting this memory module shows substantial improvements on both single and joint-memory tasks demonstrating the potential for future research in developing personalized embodied agents.
 2 Related Work
@@
 LLMs have significantly advanced embodied agents’ reasoning and planning capabilities in recent years.
-Researchers have explored LLMs for interpreting user goals [2], high-level task planning [17], and integrating LLMs into comprehensive embodied agent frameworks [18, 29, 34, 44, 19].
-Other research directions have focused on generating executable code for embodied tasks directly from language instructions [30, 47, 43], while various benchmarks have been developed to evaluate embodied reasoning abilities [26, 9, 7, 28].
+Researchers have explored LLMs for interpreting user goals (Ahn et al., 2022), high-level task planning (Huang et al., 2022a), and integrating LLMs into comprehensive embodied agent frameworks (Huang et al., 2022b; Li et al., 2022; Mu et al., 2023; Song et al., 2023; Huang et al., 2023b).
+Other research directions have focused on generating executable code for embodied tasks directly from language instructions (Liang et al., 2022; Wang et al., 2023a; Singh et al., 2023), while various benchmarks have been developed to evaluate embodied reasoning abilities (Li et al., 2023; Choi et al., 2024; Chang et al., 2024; Li et al., 2025b).
 Collectively, these studies highlight the promise of LLM-powered agents in bridging language understanding and physical interaction.
 Memory systems for embodied agents.
-Previous studies on memory systems for embodied agents have primarily focused on semantic memory (e.g., scene graph, semantic map), which store and provide state information about the current environment [39, 23, 16, 51, 48], or on procedural memory (e.g., skill library) that stores action primitives, focusing on how to perform tasks to enhance the efficiency on generating the low-level action code [47, 42, 59].
+Previous studies on memory systems for embodied agents have primarily focused on semantic memory (e.g., scene graph, semantic map), which store and provide state information about the current environment (Rana et al., 2023; Kim et al., 2023; Gu et al., 2024; Xie et al., 2024; Wang et al., 2024c), or on procedural memory (e.g., skill library) that stores action primitives, focusing on how to perform tasks to enhance the efficiency in generating the low-level action code (Wang et al., 2023a; Sarch et al., 2023; Zhang et al., 2023).
 Another important category is episodic memory, which captures specific past interactions and experiences with users.
-However, prior uses of episodic memory have mostly treated it as passive task buffers [2, 43] or histories for in-context [18, 31, 44, 8], without explicitly evaluating its role in personalized task grounding or systematic memory utilization.
+However, prior uses of episodic memory have mostly treated it as passive task buffers (Ahn et al., 2022; Singh et al., 2023) or histories for in-context (Huang et al., 2022b; Liu et al., 2023; Song et al., 2023; Chen et al., 2024), without explicitly evaluating its role in personalized task grounding or systematic memory utilization.
 Personalization for embodied agents.
-The importance of personalization in robotics has long been recognized [13, 25, 10], particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users.
-Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement [22, 49], table settings [37], or personalized object navigations [11, 4].
+The importance of personalization in robotics has long been recognized (Dautenhahn, 2004; Lee et al., 2012; Clabaugh & Matarić, 2018), particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users.
+Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement (Kapelyukh & Johns, 2021; Wu et al., 2023), table settings (Puig et al., 2021), or personalized object navigations (Dai et al., 2024; Barsellotti et al., 2024).
 Recently, Xu et al.
-[52] aim to infer user preferences from a few demonstrations and adapt planning behavior accordingly.
-However, these approaches primarily focus on implicit preference adaptation or short-term reactive behaviors, without modeling user-specific knowledge in a structured manner.
-3 Preliminaries
-We formulate the object rearrangement task for LLM-powered embodied agents as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor.
-At each timestep $t$, the environment is in a state $s_{t}\in S$, and the agent receives a partial observation $w_{t}\in\Omega$ in text modality describing visible objects near its current position.
-At the beginning of each episode, the agent is given a natural language instruction $I$ (e.g., Place the mug on the table and the book on the shelf).
-The instruction is grounded into a symbolic representation of the ground-truth goal $g$, denoted $g=\{(o_{i},l_{i})\}_{i=1}^{k}$, where each pair $(o_{i},l_{i})$ are the target object $o_{i}$ (e.g., mug) that should be placed at the specified location $l_{i}$ (e.g., on the table).
-In order to execute the instruction, the agent must internally derive the goal representation $\phi(I)\longrightarrow g$, where $\phi$ denotes the instruction grounding function, to guide the policy’s decision-making.
-Given an instruction $I$, where the policy $\pi$ is implemented by an LLM, the agent generates actions at timestep $t$ based on the trajectory of observations and actions:
-$\pi(I,\tau_{t})\rightarrow a_{t},\quad\tau_{t}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{t-1},a_{t-1},w_{t})$ (1)
-The goal is to produce a sequence of actions $a_{1:t}=(a_{1},a_{2},\ldots,a_{t})$ such that the resulting state $s_{t}$ satisfies the agent’s goal $g(s_{t})=1$.
-4 Memento
-In this section, we introduce Memento, a personalized embodied agent evaluation framework designed to assess how well embodied agents leverage episodic memory containing personalized knowledge to provide personalized assistance.
-We begin by describing how we design the memory evaluation process (§ 4.1), categorize personalized knowledge (§ 4.2), the data construction process (§ 4.3), and the validation of our evaluation framework (§ 4.4).
-Figure 2: Overview of Memento.
-4.1 Memory Evaluation Process Design
-Two-stage evaluation process.
-The major challenge of evaluating embodied agents’ memory utilization capability is to quantify the memory effect on the overall task performance.
-To address this limitation, as shown in Figure 2, we divide the evaluation process into two stages.
-  • Memory Acquisition Stage: Agents perform tasks with conventional object rearrangement task instructions containing personalized knowledge while accumulating the interaction history (i.e., episodic memory).
-The goal of this stage is to provide a reference performance baseline.
-  • Memory Utilization Stage: Agents execute the same tasks as in the memory acquisition stage but with modified instructions that require agents to recall and apply the previously acquired personalized knowledge to succeed.
-The goal of this stage is to evaluate how well agents can utilize memory by comparing the performance drop relative to the acquisition stage.
+(2024) aim to infer user preferences from a few demonstrations and adapt planning behavior accordingly.
+However, these approaches primarily address implicit preference adaptation or short-term reactive behaviors derived from repeated user demonstrations, without leveraging explicitly provided, user-defined knowledge in a structured manner.
+3 Evaluation Setup
+3.1 Task & Focus
+Task: personalized object rearrangement.
+We formulate the personalized object rearrangement task as a Partially Observable Markov Decision Process (POMDP) $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor.
+At timestep $t$, the agent receives a text observation $w_{t}\in\Omega$ describing nearby objects and generates actions via LLM policy $\pi$.
+Given a natural language instruction $I$ that may contain personalized references (e.g., "Place my favorite mug on the table"), the agent must derive the goal representation $g$ using both the instruction and personalized knowledge from memory: $\phi(I,M)\longrightarrow g={(o_{i},l_{i})}_{i=1}^{k}$, where $M={h_{1},h_{2},\ldots,h_{n}}$ represents the set of episodic memories from previous interactions, and each $(o_{i},l_{i})$ pair specifies target object $o_{i}$ and location $l_{i}$.
+The policy generates actions based on the observation-action trajectory $\tau_{t}$:
+$\pi(I,M,\tau_{t})\rightarrow a_{t},\quad\tau_{t}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{t-1},a_{t-1},w_{t})$ (1)
+The goal is to produce action sequence $a_{1:t}$ such that the resulting state $s_{t}$ satisfies all object-location specifications in $g$.
+Focus: memory utilization.
+Personalized assistance tasks require agents to integrate multiple capabilities including natural language understanding, memory retrieval, planning, and action execution.
+Among these interconnected abilities, we focus specifically on memory utilization—the agent’s capacity to effectively recall and apply relevant personalized knowledge from episodic memory $M$ to interpret instructions and guide task planning.
+This focus is motivated by memory utilization being the primary distinguishing capability from an agent perspective that separates personalized assistance tasks from conventional embodied tasks, making it essential to investigate for potential challenges and solutions for developing personalized embodied agents.
+Figure 2: Overview of Memento. The framework evaluates memory utilization capability by comparing agent performance on tasks with identical goals but varying instructions on each stage.
+3.2 Dimensions of Personalized Knowledge Analysis
+We analyze embodied agents’ memory utilization capabilities along two dimensions that reflect how personalized knowledge affects goal interpretation in $\phi(I,M)\rightarrow g={(o_{i},l_{i})}_{i=1}^{k}$.
+Therefore, we categorize personalized knowledge into two types, each designed to isolate distinct reasoning challenges that agents must resolve using episodic memory:
+  • Object semantics evaluates agents’ ability to identify target objects $o_{i}$ based on personal meaning assigned by users, including ownership (my cup), preference (my favorite running gear), personal history (a graduation gift from my grandma), or grouped references (my childhood toy collections).
+This dimension tests whether agents can recall specific object semantics from prior interactions to resolve ambiguous object references.
+  • User patterns assesses agents’ capacity to reconstruct complete goals $g$ by leveraging consistent behavioral patterns, including personal routines (my remote work setup) and arrangement preferences (my cozy dinner atmosphere).
+This dimension evaluates whether agents can apply previously observed patterns across multiple objects and locations to infer task requirements.
+Each dimension targets different aspects of the grounding function $\phi(I,M)\rightarrow g$: object semantics focuses on resolving individual object references, while user patterns addresses holistic goal reconstruction from behavioral context.
+Detailed explanation of the sub-categories of personalized knowledge can be found in Appendix C.2.
+3.3 Memento: Personalized Embodied Agent Evaluation Framework
+Framework design.
+The major challenge of evaluating embodied agents’ memory utilization capability is quantifying the memory effect on overall task performance.
+To address this, as shown in Figure 2, we design an end-to-end two-stage evaluation process that measures memory utilization capabilities through instruction interpretation.
 Given the base object rearrangement task episode defined as the tuple $\epsilon=(S,I,g)$, the key concept of our evaluation process is to share the scene $S$ and goal representation $g$, while varying the instruction $I$ across the two stages to isolate instruction interpretation capability as the primary factor influencing performance differences.
-Formally, in the memory acquisition stage, each episode is defined as $\epsilon_{acq}=(S,I_{acq},g)$, where the instruction $I_{acq}$ contains sufficient information to infer the goal $g$, denoted as:
-$\phi(I_{acq})\longrightarrow g$ (2)
-During this stage, we also store the episodic memory $h_{acq}$:
-$h_{acq}=(I_{acq},\tau_{k})\in H_{acq},\quad\tau_{k}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{k-1},a_{k-1},w_{k})$ (3)
-In the subsequent memory utilization stage, each episode is defined as $\epsilon_{util}=(S,I_{util},g)$, where the instruction $I_{util}$ is intentionally underspecified and requires the agent to recall the corresponding episodic memory $h_{acq}$ to correctly interpret the goal $g$, formally:
-$\phi(I_{util},h_{acq})\longrightarrow g$ (4)
-Through this design concept, we are able to quantify the agent’s ability to utilize memory by comparing performance between the two stages.
-Memory utilization stage task design.
-Within our two-stage evaluation process design, we can assess embodied agents’ ability to utilize personalized knowledge from a single episodic memory.
-However, this approach alone fails to capture real-world complexity and lacks analytical diversity for comprehensive assessment.
-Therefore, to evaluate different levels of memory complexity, we divide our assessment into (1) Single-memory task, which require utilizing information from one episodic memory, and (2) Joint-memory task, which necessitate synthesizing information from two distinct episodic memories to successfully complete the episode.
-Building on our evaluation process concept, we form the joint-memory task by concatenating two episodes, which can be formulated as:
-$\epsilon_{util}^{joint}=(S,I_{util}^{joint},[g^{i};g^{j}])$ (5)
-where $i$, $j$ denotes the corresponding reference episodes from the memory acquisition stage.
-4.2 Personalized Knowledge Categorization
-Building upon our memory evaluation process, we aim to analyze embodied agents’ ability to understand personalized knowledge in object rearrangement tasks by focusing on its role in goal interpretation, $\phi(I)\rightarrow g=\{(o_{i},l_{i})\}_{i=1}^{k}$, where each $o_{i}$ and $l_{i}$ denote the target object and location, respectively.
-To facilitate this analysis, we categorize personalized knowledge into two types, each comprising subcategories  [^1] A detailed explanation of the sub-categories of personalized knowledge can be found in Appendix C.1.
-that reflect how users naturally express preferences and routines in real-world interactions.
-Each type is designed to isolate a distinct reasoning challenge that the agent must resolve by utilizing episodic memory during the memory utilization stage.
-  • Object semantics: Individual objects that the user assigns personal meaning, encompassing subcategories such as ownership (e.g., my cup), preference (e.g., my favorite running gear), past history (e.g., a graduation gift from my grandma), or grouped references (e.g., my childhood toy collections).
-This category tests whether the agent can identify the target object $o_{i}$ by recalling its personal meaning from prior interactions.
-  • User patterns: Sequences of actions that the user consistently performs, including personal routines (e.g., my remote work setup) and arrangement preferences (e.g., my cozy dinner atmosphere) across recurring contexts.
-This category evaluates the agent’s ability to reconstruct the complete goal $g$ by leveraging previously observed behavioral patterns across multiple objects and locations.
-4.3 Dataset Construction Process
-We constructed the dataset for Memento through a four-step process using the Habitat 3.0 simulator [38] as the environment, with a simulated Spot robot as the agent [5, 58].
-Our custom dataset spans 12 scenes, comprising a total of 438 episodes distributed across stages.
-Notably, the memory acquisition stage and the single-memory task in the utilization stage have the same number of episodes, whereas the joint-memory task contains fewer episodes.
-The detailed dataset statistics and the explanation of the construct process is described in Appendix C.3, and Appendix C.4.
-Step 1: Object rearrangement task collection.
-We use the test set of the PartNR [7] as our foundation object rearrangement task data.
-Unlike simple pick-and-place tasks, PartNR episodes require completing multiple object–location pairs within a single instruction, which aligns with our memory evaluation process design.
-Step 2: Scene augmentation with distractor objects.
-In the original scenes, there was an issue where no objects of the same type as the target object were present.
-As a result, the agent could identify the target object without needing to understand personalized knowledge.
-To address this, we augmented the scenes by placing distractor objects of the same type near the target object.
-For example, if the target object is a “blue cup” on the table, we place a “red cup” next to it as a distractor.
-Step 3: Task instruction generation.
-We first generate personalized knowledge contextually tailored to the original task instruction using GPT-4o.
-With the generated personalized knowledge, we applied to both stage instruction curation.
-As illustrated in Figure 2, the memory acquisition stage instruction $I_{acq}$ is constructed by concatenating the base instruction, object visual captions (only for episodes of the object semantics type), and the generated personalized knowledge.
-This ensures that the goal can be inferred directly from the instruction.
-For the memory utilization stage, we prompt GPT-4o to generate a personalized instruction that implicitly reflects the personalized knowledge, based on the same base instruction.
-For joint-memory tasks, we concatenate two such personalized instructions sequentially.
-Figure 3: The performance results without using episodic memory. Original indicates the conventional object rearrangement task episodes.
-Step 4: Quality control.
-To ensure data quality, we first heuristically filtered episodes containing similar memories referencing identical objects within scenes, preventing interference between similar episodic memories.
-Subsequently, we manually reviewed episodes from the memory acquisition stage where GPT-4o failed to successfully complete the task, where we filtered out episodes that contained unnatural instructions or cases where the generated instructions did not match the intended goal representation, ensuring the quality of our evaluation data.
-4.4 Validation of Memento
-To validate that Memento effectively assesses embodied agents’ memory utilization capability, we compare performance across stages in a setup without memory retrieval.
-As shown in Figure 3, compared to the results of the original object rearrangement task and the memory acquisition stage, embodied agents struggle to complete tasks in the memory utilization stage.
-Since the underlying episode remains the same across all tasks, this result confirms that the instruction $I_{util}$ in the memory utilization stage is difficult to interpret without access to previous interaction histories.
-Notably, we observed a particular behavior of the embodied agents: when there are two objects of the same type, agents that do not understand which one is the actual target tend to randomly select one and proceed with the task.
-This explains the reason why performance in the single-memory task is higher than that of the joint-memory task.
-5 Evaluating Personalized Knowledge Utilization in Episodic Memory
-5.1 Experimental Setup
-Evaluation metrics.
-Following Chang et al.
-[7], we use two main metrics: Percent Complete (PC) for the proportion of goal completion, and Success Rate (SR) for full task completion.
-We also report Sim Steps, which show the number of simulation steps required for agents to complete the task, and Planning Cycles, which indicate the number of LLM inference calls made during task execution.
-To evaluate memory utilization, we also report performance drops between acquisition and utilization stages as $\Delta$PC and $\Delta$SR.
-Note that for joint-memory tasks, these differences are computed relative to the average performance of the corresponding acquisition-stage episodes.
-Implementations.
-Following prior work [45, 38, 7], we implement a LLM-powered embodied agent architecture, where the LLM functions as a high-level policy planner that selects appropriate skills from a predefined skill library.
-We use ReAct [56] prompt format for LLMs to take actions.
-Additionally, we implement a top-5 memory retrieval setup  [^2] We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer [40].
-for the memory utilization stage, ensuring the corresponding memory is included in the retrieved results by randomly replacing one memory if the correct one is not initially retrieved.
-Full implementation details are in AppendixB.2.
-Models.
-We evaluate embodied agents powered by a range of LLMs to compare memory utilization capabilities across model families and sizes, including proprietary models (GPT-4o [21], Claude-3.5-Sonnet [3]) and open-source models (Llama-3.1-70b/8b [15], Qwen-2.5-72b/7b [54]).
+In the memory acquisition stage, agents perform conventional object rearrangement tasks with instructions that contain sufficient personalized knowledge.
+These episodes are defined as $\epsilon_{acq}=(S,I_{acq},g)$ where $\phi(I_{acq})\longrightarrow g$, meaning agents can directly infer the goal from instructions.
+During this stage, agents establish a reference performance baseline while accumulating episodic memory $h_{acq}$ from their interactions.
+The memory utilization stage presents agents with identical tasks but uses underspecified instructions $\epsilon_{util}=(S,I_{util},g)$.
+These tasks can only be completed when agents recall and apply relevant personalized knowledge, requiring the extended grounding function $\phi(I_{util},h_{acq})\longrightarrow g$.
+Through this evaluation process, we are able to quantify the agent’s memory utilization capability by comparing performance between the two stages.
+Furthermore, to evaluate different complexity levels, we assess both (1) single-memory tasks requiring information from one episodic memory, and (2) joint-memory tasks necessitating synthesis from two distinct memories, formulated as $\epsilon_{util}^{joint}=(S,I_{util}^{joint},[g^{i};g^{j}])$ where $i,j$ denote corresponding reference episodes from the acquisition stage.
+As shown in Figure 3, we conduct experiments without memory usage to validate our framework design, confirming that agents struggle significantly with underspecified instructions.
+Figure 3: The performance results without using episodic memory. Original indicates episodes from PartNR dataset.
+Dataset construction.
+We construct the dataset for Memento using Habitat 3.0 simulator (Puig et al., 2023) with a simulated Spot robot (Boston Dynamics, 2025; Yokoyama et al., 2023).
+Our dataset spans 12 scenes with a total of 438 episodes, using PartNR (Chang et al., 2024) test set as foundation data since its multi-object rearrangement tasks align with our evaluation design.
+Using GPT-4o, we generate personalized knowledge and create two instruction types: $I_{acq}$ concatenating base instruction, visual captions, and personalized knowledge for direct goal inference; and $I_{util}$ containing personalized references requiring memory recall.
+To prevent agents from identifying targets without personalized knowledge, we augmented scenes by placing distractors of the same type but with different appearances near target objects (e.g., placing a "red cup" next to the target "blue cup").
+For quality control, we filtered episodes with similar memories referencing identical objects and manually reviewed failed cases to ensure instruction quality.
+We further validate the dataset with an inter-rater reliability analysis, achieving an average Krippendorff’s alpha of 0.69 across four key criteria.
+Details of the Memento are provided in Appendix C.3.
+3.4 Metrics & Retrieval Setup
+[…]
+Retrieval setup.
+For memory retrieval in the utilization stage, we use top-$k$ similarity-based retrieval with all-mpnet-base-v2 (Reimers & Gurevych, 2019), treating the current instruction as a query and episodic memory instructions as keys.
+Since our primary objective is evaluating memory utilization capability rather than retrieval performance, we ensure the corresponding gold memories are always included in the top-$k$ results through randomly replacing retrieved candidates when necessary.
+We set $k=5$ as our default configuration because this achieves 96.5% recall for gold memories in realistic retrieval scenarios, compared to only 86.1% recall at $k=3$.
+4 Evaluating Memory Utilization Across Embodied Agents
+Through Memento, we assess memory utilization capabilities across multiple LLM-powered embodied agents of varying parameter sizes to understand how effectively they provide personalized assistance and identify their limitations (RQ1).
+4.1 Models & Implementation Details
+We evaluate both proprietary models (GPT-4o (Hurst et al., 2024), Claude-3.5-Sonnet (Anthropic, 2024)) and open-source models (Llama-3.1-70b/8b (Grattafiori et al., 2024), Qwen-2.5-72b/7b (Yang et al., 2024)) with different model families and sizes.
+We implement LLM-powered embodied agents following established architectures (Szot et al., 2021; Puig et al., 2023; Chang et al., 2024), where LLM serves as a high-level policy planner that selects appropriate skills from a predefined skill library using ReAct (Yao et al., 2023) prompting format.
+Note that, we use gold perception and motor skills to focus on challenges related to embodied agents’ cognitive abilities.
+Complete implementation details are provided in Appendix C.1.
 Table 1: Model performance across memory acquisition and utilization stage in Memento.
-Model Stage Task Type (Memory) Planning Cycles $\downarrow$ Sim Steps $\downarrow$ Percent Complete $\uparrow$ $\Delta$ PC Success Rate $\uparrow$ $\Delta$ SR
+Model Stage Task Type (Memory) Planning Cycles $\downarrow$ Sim Steps $\downarrow$ Percent Complete $\uparrow$ $\Delta$ PC $\uparrow$ Success Rate $\uparrow$ $\Delta$ SR $\uparrow$
 GPT-4o Acquisition - 16.5 2156.1 96.3 - 95.0 -
@@
 Joint 26.9 4149.0 33.7 -34.2 5.6 -52.7
-5.2 Main Results
+4.2 Results
+Figure 4: The results of personalized knowledge type based analysis (single-memory).
 LLM-powered embodied agents struggle with understanding personalized knowledge.
-As shown in Table 1, while GPT-4o maintains a relatively high success rate in the single-memory task, all models show a success rate drop over 20% compared to the memory acquisition stage.
-In particular, for joint-memory tasks even GPT-4o exhibits a 30.5% drop in success rate, highlighting the increased difficulty of these settings.
-This substantial performance decline demonstrates that even frontier models struggle to accurately reference personalized knowledge from episodic memory, and often fail to consistently apply it across multiple steps in long-horizon task planning.
-Figure 4: The results of personalized knowledge type based analysis (single-memory).
-LLM-powered embodied agents exhibit increased exploration behavior on joint-memory tasks.
-Joint-memory task results reveal that LLMs (even GPT-4o) find it difficult to recall personalized knowledge from different memory sources.
-The number of planning cycles and simulation steps significantly increases compared to other tasks, this suggests that the embodied agent fails to correctly interpret the instruction, leading to excessive exploration during task execution.
-Also the performance gap between percent complete and success rate is larger than in the single-memory task, which indicates that the agent frequently misses part of the necessary information for successful task completion.
-5.3 In-depth Analysis
-To better understand the challenges of providing personalized assistance, we examine two key aspects: the root causes of their limitations and how the top-$k$ memory setting affects memory utilization.
-5.3.1 Personalized Knowledge Type-based Analysis
-We analyze the performance gap between the memory acquisition stage and the single-memory task from the memory utilization stage by comparing success rates across different types of personalized knowledge.
-Analysis results for the joint-memory task are provided in Appendix D.2.
-LLMs can recall objects but struggle to comprehend action sequences.
-As shown in Figure 4, all models exhibit substantially larger performance drops for tasks requiring understanding user patterns compared to object semantics.
-Notably, the performance drop in object semantics is minimal, indicating that LLMs are relatively effective at directly recalling relevant memories to identify target objects.
-In contrast, tasks involving user patterns pose a significantly greater challenge, as they require integrating and reasoning over sequences of events.
-5.3.2 Memory Top-K Analysis
-We analyze how varying the number of retrieved memories ($k$) impacts the agent’s ability to utilize personalized knowledge from episodic memory.
-Irrelevant memories act as a distraction to agents.
-In Figure 5, we observe that as $k$ increases, all models exhibit consistent performance degradation across both task types, highlighting the difficulty of identifying the exact information from a growing set of retrieved memories.
-Consistent with previous findings, the degradation is especially pronounced in tasks requiring user patterns understanding, suggesting that such tasks are particularly vulnerable to noise when recalling and executing multi-step procedures grounded in implicit, personalized knowledge.
+In Table 1, we observe that while GPT-4o maintains a relatively high success rate in the single-memory tasks, all models show a success rate drop over 20% compared to the memory acquisition stage.
+This substantial performance decline across all evaluated models demonstrates that current embodied agents cannot reliably integrate personalized knowledge for task execution.
+Current embodied agents can effectively recall object semantics but struggle to comprehend user patterns.
+We then analyze the agents performance on single-memory task by personalized knowledge types.
+Notably, as shown in Figure 4, all models show minimal performance degradation for object semantics tasks, while user pattern tasks exhibit substantial performance drops across all evaluated models, indicating that current embodied agents face significant limitations in comprehending sequential information within behavioral contexts.
+Figure 5: Taxonomy of successful and failed cases in memory utilization with illustrative examples.
+Top: success and failure cases of object semantics; Bottom: success and failure cases of user patterns.
+Agents default to commonsense over personalized knowledge.
+To understand the agents’ behavior on personalized assistance tasks, we conduct qualitative case study which is shown in Figure 5.
+Through this analysis, we identify an interesting pattern: agents tend to rely on commonsense knowledge rather than personalized memory.
+For example, when reasoning about "afternoon refresh setup" (C.1), the agent relies on commonsense knowledge rather than referencing personalized memory, correctly predicting "jug and cup" because this commonsense happens to align with the user’s actual behavior.
+However, when applying the same commonsense-based reasoning approach to "playtime" (D.1), the agent incorrectly places toys on the TV room table instead of consulting the personalized memory that specifies the user’s preference for placing toys on the bench.
+This indicates that when agents struggle to reference personalized knowledge from memory, they default to parametric commonsense knowledge rather than leveraging non-parametric personalized knowledge, leading to inconsistent performance across user contexts.
+5 Scenario-based Analysis for Memory Utilization Bottlenecks
+In this section, we conduct scenario-based analysis to understand the factors that negatively affect agents’ memory utilization capabilities in personalized assistance tasks (RQ2).
+First, we examine agent performance when increasing the number of retrieved memories in the context (§˜5.1).
+Second, we analyze agent behavior when tasks require coordinating multiple memories to achieve task completion (§˜5.2).
 Figure 6: Success rate comparison across models as top-k value increases. Dashed lines represent memory acquisition stage baselines for each model.
-Figure 6: Examples of success and error cases in memory utilization stage. Top: success and failure cases of object semantics; Bottom: success and failure cases of user patterns.
-5.4 Empirical Analysis of Success and Error Cases
-To better understand how LLM-powered embodied agents utilize personalized knowledge derived from episodic memory during task planning, we conduct qualitative case studies of both successful and failed episodes.
-Figure 6 presents the taxonomy of case types along with illustrative examples.
-Detailed explanations for each case type are provided in Appendix D.4.
-Below, we highlight a set of representative cases that offer key insights into the embodied agents’ behavior.
-  • Agent misses personalization cues (B.1): The agent fails to recognize user-specific references and treats them as generic or proper nouns, stemming from incorrect interpretations of user intent.
-  • Use of commonsense knowledge over personalized knowledge (C.1 & D.1): Even when relevant episodic memory is available, the agent often relies on general commonsense knowledge to infer user routines, rather than using the personalized knowledge.
-This tendency appears in both successful and failed episodes.
-6 Discussion
-6.1 The Impact of Trajectory in Episodic Memory
-Table 2: Analysis of how memory type affects agent performance: (a) complete action action trajectories containing user instructions, (b) summary of (a), and (c) user instructions only.
+5.1 Analyzing Agent Performance Under Increased Memory Context
+To understand whether agents remain robust in identifying correct information as memory volume increases, we vary the number of retrieved memories (top-$k$) in single-memory tasks.
+We ensure gold memory is always included and focus on whether agents can reference the correct memory.
+Information overload degrades memory utilization performance.
+As shown in Figure 6, increasing $k$ consistently degrades performance across all models and both personalized knowledge types, indicating that agents struggle to extract relevant information when presented with larger sets of retrieved memories.
+We found that agents increasingly rely on commonsense knowledge rather than personalized memories as retrieved memories increase, especially on user patterns, suggesting that embodied agents suffer from information overload.
+Therefore, even though LLMs demonstrate impressive capabilities in processing long-context information (Wang et al., 2024b; Liu et al., 2025), memory systems that retrieve appropriate numbers of memories are crucial for personalized embodied agents.
+We provide further analysis of retrieval strategies in scenarios without guaranteed gold memory access in Appendix E.3.
+5.2 Investigating Agent Behavior in Joint-memory Tasks
+In realistic scenarios, users often refer to multiple pieces of personalized knowledge within a single instruction.
+Accordingly, we fix the top-$k$ parameter to $k=5$ and employ joint-memory tasks in Memento to analyze the agents’ performance on tasks that require synthesizing information from two distinct episodic memories.
+Figure 7: Personalized knowledge type-based analysis on joint-memory tasks.
+[…]
+6 Exploring Memory Design for Personalized Embodied Agents
+Table 2: Model performance with memory format simplification approaches. Full memory indicates the episodic memory.
 Model Memory PC (%) SR (%)
-GPT-4o (a) 90.0 83.3
-(b) 88.0 83.3
-(c) 62.4 50.0
-Qwen-2.5-72b (a) 77.2 66.7
-(b) 77.4 70.0
-(c) 51.3 40.0
-Llama-3.1-8b (a) 72.8 63.3
-(b) 49.4 43.3
-(c) 40.0 30.0
-Qwen-2.5-7b (a) 50.1 43.3
-(b) 43.9 36.7
-(c) 35.6 23.3
-Memento’s design incorporates complete action-observation trajectories in episodic memory, raising the question of whether agents should reference these detailed trajectories rather than relying solely on user instructions.
-To investigate this, we conduct an additional comparative experiment to evaluate if providing full action trajectories offers significant advantages over simply providing instructions that state personalized knowledge.
-For the experiment setup, we evaluate model performance across three different cases of memory provided to the agent as shown in Table 2.
-The results reveal that while larger models (GPT-4o, Qwen-2.5-72b) perform well with only high-level plans (b), smaller models (Llama-3.1-8b, Qwen-2.5-7b) require full procedural details from completed trajectories (a) to succeed.
-Most notably, all models show substantial performance drops when given only user instructions (c), suggesting that action trajectories contain essential procedural cues necessary for understanding personalized knowledge, regardless of model capacity.
-Experiment setup details are provided at Appendix D.5.
-6.2 Agent Behavior under Ambiguous Instructions from Users
-Table 3: Performance under ambiguous queries for personalized knowledge.
-PC (Percent Complete) and SR (Success Rate) (%) indicate how well agents resolve indirect references to personalized knowledge from memory.
-Model PC (%) SR (%)
-GPT-4o (Baseline) 92.0 90.0
-GPT-4o 80.4 73.3
-Qwen-2.5-72b (Baseline) 75.1 66.7
-Qwen-2.5-72b 59.6 53.3
-While Memento focuses on evaluating personalized knowledge grounding with explicit references to prior interactions, real-world human-agent communication often involves ambiguous or indirect references.
-To explore this challenge, we conducted a proof-of-concept experiment to assess whether current models can interpret ambiguous instructions that indirectly refer to previously encountered personalized knowledge.
-[^3] Further details of the dataset and experiment setup are provided in Appendix D.5.
-We created a tailored set of tasks referencing personalized knowledge using contextual cues, synonyms or causal references (e.g., Can you set my afternoon tea time routine?
-$\rightarrow$ I’m about to enjoy my afternoon tea.
-Could you set things up as I like them?).
-The results, shown in Table 3, reveal the degradation in performance, indicating that handling ambiguous queries remains a key challenge for future personalized embodied agents.
-We view this as a promising direction for future work, where we plan to systematically extend Memento to evaluate LLM-powered embodied agents’ capabilities under ambiguous and implicit reference scenarios, aiming to better reflect the complexity of real-world human-agent interaction.
+GPT-4o Full Memory 90.0 83.3
+Summarization 88.0 83.3
+Instruction-only 62.4 50.0
+Qwen-2.5-72b Full Memory 77.2 66.7
+Summarization 77.4 70.0
+Instruction-only 51.3 40.0
+Llama-3.1-8b Full Memory 72.8 63.3
+Summarization 49.4 43.3
+Instruction-only 40.0 30.0
+Qwen-2.5-7b Full Memory 50.1 43.3
+Summarization 43.9 36.7
+Instruction-only 35.6 23.3
+Based on our previous findings that demonstrate critical bottlenecks in memory utilization for personalized assistance, we identify that these limitations stem from the fundamental challenge of extracting relevant personalized knowledge from episodic memories.
+We hypothesize that clearly providing personalized knowledge information—an approach consistent with recent advances in context engineering (Mei et al., 2025)—will enhance agents’ ability to leverage memory effectively.
+Therefore, in this section, we explore memory architectural approaches to improve agents’ memory utilization capabilities for personalized assistance tasks (RQ3).
+6.1 Investigating Memory Format Simplification
+As episodic memory contains full interaction histories, it includes contextual information irrelevant to personalized knowledge extraction.
+Therefore, we investigate memory format simplification approaches that provide agents with refined information by removing irrelevant content, hypothesizing that this reduction of noise will improve performance.
+We compare full episodic memory against two simplified alternatives: (1) GPT-4o summarized episodic memory that retains key information while removing trajectory details, and (2) instruction-only setups that rely solely on task instructions, since personalized knowledge can be extracted from instructions alone.
+Episodic memory provides both personalized knowledge and in-context learning benefits.
+As shown in Table 2, summarized memory shows minimal impact on large models (GPT-4o, Qwen-2.5-72b).
+Notably, smaller models (Llama-3.1-8b, Qwen-2.5-7b) exhibit performance degradation when using summarized memory compared to full episodic memory.
+This performance drop occurs because summarizing removes the trajectory information that provides crucial in-context learning benefits for LLMs, reducing the overall task success rate.
+These findings demonstrate that episodic memory is an essential module for embodied agents.
+6.2 Developing Hierarchical Knowledge Graph-based User Profile Memory
+Figure 8: Agent performance with user profile memory across single- and joint-memory tasks.
+Building upon our finding that episodic memory is an essential memory module, we devise a hierarchical knowledge graph-based user profile memory module that manages personalized knowledge independently to provide cleaner, more accessible personalized information to the agent.
+Figure 9: Illustration of user profile memory.
+Memory module design.
+A key challenge in developing a personalized knowledge module is organizing knowledge relationships effectively.
+Sequential knowledge like user patterns requires sophisticated memory structure to prevent information loss or failing to adapt to evolving knowledge (e.g., "Add the mug from my coffee set to my morning routine and prepare my routine").
+Inspired by previous works on knowledge graphs (Speer et al., 2017; Bosselut et al., 2019), we construct personalized knowledge as a hierarchical knowledge graph that captures both object semantics and user patterns.
+As shown in Figure 9, user profile memory employs a three-level structure with users at the top, knowledge types (object semantics and user patterns) in the middle, and specific elements (objects, patterns, locations) at the bottom.
+The graph connects these levels using hierarchical edges for structural relationships and temporal edges for sequential ordering within user patterns.
+This design enables systematic representation and easy updates of personalized knowledge—for instance, a red mug from a coffee set can be dynamically added to a morning routine by updating the relevant knowledge relationships.
+More details of the user profile memory design, schema, extraction and retrieval algorithms are provided in Appendix D.1.
+User profile memory enhances embodied agents’ memory utilization capability.
+As shown in Figure 8, agents with user profile memory achieves significant performance improvements across all models for both single-memory and joint-memory tasks.
+The performance gains are especially notable on user pattern tasks, where the sequential structure had previously posed challenges.
+These results indicate that user profile memory effectively manages personalized knowledge, demonstrating its ability to enhance agents’ memory utilization capabilities for personalized assistance tasks.
 7 Conclusion
-In this work, we present Memento, a novel personalized embodied agent evaluation framework designed to assess the LLM-powered embodied agents’ ability to utilize episodic memory for providing personalized assistance.
-Our experiments across a range of LLM-powered embodied agents reveal key limitations in their ability to effectively leverage personalized knowledge from memory, particularly when integrating multiple memories and interpreting user patterns.
-These findings highlight the gap between current capabilities and the demands of real-world personalized assistance.
-We hope that Memento serves as a stepping stone for future research in developing more effective personalized embodied agents.
-Appendix A Limitations
+This work investigates memory utilization challenges in LLM-powered embodied agents for personalized assistance through Memento, revealing that while agents can recall simple object semantics, they struggle with sequential user patterns, suffer from information overload when processing multiple retrieved memories, and fail to coordinate joint-memory tasks effectively.
+We demonstrate that episodic memory provides both personalized knowledge and in-context learning benefits, explaining why simple summarization approaches fail.
+To address these bottlenecks, we design a hierarchical knowledge graph-based user-profile memory module that separately manages personalized knowledge while preserving episodic memory benefits, achieving substantial improvements across both single-memory and joint-memory tasks.
+We hope these findings will inspire future research toward more sophisticated memory architectures that enable truly effective personalized embodied agents.
+Ethics Statement
+This work involves no human subjects, conflicts of interest, or legal compliance concerns.
+We use licensed resources (e.g., PARTNR (Chang et al., 2024) under the MIT License) and commonsense knowledge without personal data, copyrighted content, or bias-amplifying material.
+We generated personalized knowledge from LLM-synthesized descriptions with prompts designed to reduce biased or discriminatory content, and we manually verified that the resulting user knowledge was free of such concerns.
+We acknowledge that embodied memory systems may pose privacy and security risks in real-world deployment.
+To mitigate these issues, future systems should adopt on-device processing, minimize retention, require user consent, and apply secure, transparent memory handling.
+We will release only de-identified or synthetic assets and provide documentation of safeguards to reduce misuse risk.
+Reproducibility Statement
+To facilitate full reproducibility of our results, we include complete evaluation dataset details and experimental procedures in the appendices.
+Furthermore, we publish our code at https://anonymous.4open.science/r/MEMENTO with thorough documentation.
+Appendix A Limitations, Privacy, Societal Impacts and LLM Usage
 A.1 Limitations
-Controlled simulator environment.
-Our experiments are conducted entirely in a controlled simulator environment [38], which does not fully reflect the complexities of real-world robotics such as perception noise, actuation uncertainty, or unstructured environments.
-Visual perception is not addressed.
-Our framework deliberately isolates memory-centric planning by excluding visual perception components such as VLMs [55] or VLA [6, 24] models.
-This enables focused evaluation of memory utilization, but limits the system’s generalizability to fully grounded perception scenarios.
-Oracle skills.
-We intentionally employ oracle skills for low-level perception and motor control to isolate and focus on the high-level planning and memory reasoning capabilities of the LLM-based agents.
-As a result, our framework does not evaluate the embodied agent’s full task execution process.
-A.2 Societal Impacts
+This work focuses on agents’ memory utilization capabilities for providing personalized assistance.
+While our study provides valuable insights, several limitations must be acknowledged.
+First, our experiments were conducted in a controlled simulator environment (Puig et al., 2023).
+Although we focused on agents’ cognitive abilities and expect marginal impact on our results, this controlled environment may not fully capture the complexity and variability of real-world, presenting potential risks for generalization.
+Second, we utilize oracle perception and motor skills to isolate memory utilization as the primary performance factor.
+Recent research has increasingly incorporated visual components into embodied agents (Huang et al., 2023a; Yang et al., 2025) or developed Vision-Language-Action (VLA) models that generate end-to-end actions (Brohan et al., 2023; Kim et al., 2024; Bjorck et al., 2025).
+Additionally, recent studies have shown that primary limitations in embodied tasks often exist at the observation and action levels (Chang et al., 2024; Sapkota et al., 2025).
+Therefore, while our focus on language-based memory utilization capabilities provides clear insights into this specific aspect, it may not capture all failure modes present in actual embodied agents.
+Finally, our framework lacks conversational components that are crucial for effective personalized assistance.
+Beyond memory utilization, personalized embodied agents require robust communication capabilities, including the ability to clarify ambiguous instructions, proactively engage with users, and adapt their communication style to individual preferences.
+Investigating how memory utilization integrates with conversational abilities to enable more comprehensive personalized assistance represents an important direction for our ongoing research.
+A.2 Privacy and Security Considerations
+As our work involves personalized knowledge and agent memory systems, we recognize potential privacy risks including unauthorized access to user-specific knowledge and inference of sensitive personal information from interaction patterns.
+While our current study utilizes synthetic personalized data to mitigate privacy concerns, real-world deployment of personalized embodied agents requires robust privacy safeguards including encryption of user-specific knowledge bases, secure access controls with proper authentication protocols, data minimization strategies, and mechanisms for user control over personal data.
+The specific implementation of these privacy protection mechanisms should depend on deployment context and applicable regulatory requirements, and we recommend thorough privacy impact assessments before real-world deployment.
+A.3 Societal Impacts
 Our work explores how embodied agents can remember and adapt to user-specific preferences through episodic memory, enabling more personalized and natural interactions.
 This capability has the potential to enhance convenience, efficiency, and user satisfaction in everyday environments—benefiting a wide range of users regardless of age or ability.
-However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents [41, 25].
+However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents (Rueben & Smart, 2016; Lee et al., 2012).
 Since episodic memory involves storing interaction history, future systems must consider secure and transparent memory handling.
@@
 We hope that Memento serves as a foundation for future research in building safe, privacy-aware, and user-aligned personalized embodied agents.
-Appendix B Details of Experiment Setup
-B.1 Evaluation Method
-We adopt the official evaluation protocol from the PartNR benchmark [7], which provides a python-based framework for assessing multi-step rearrangement tasks.
-We use this framework without modification.
-The evaluator analyzes simulator states using three components: (1) propositions that define object relationships to satisfy (e.g., is_on_top([spoon_1], [table_1])), (2) dependencies that define temporal conditions for multi-step instructions (after_satisfied, after_unsatisfied), and (3) constraints that enforce execution requirements (e.g., ordering, object consistency across steps).
-We rely on this system to evaluate tasks involving ambiguous references and sequential dependencies.
-The evaluation produces a percent-complete score and binary success indicator.
-B.2 LLM-Powered Embodied Agent Architecture
+A.4 LLM Usage
+Our use of LLMs was for the following specific purposes:
+  • Writing and Editing Assistance: We utilize LLMs for improving the clarity, grammar, and readability of the text.
+This included correcting grammatical errors, and ensuring consistent terminology throughout the paper.
+  • Code Generation for Experiments: We utilize LLMs to accelerate the implementation of our experimental framework.
+The authors meticulously reviewed, edited, and validated all content generated or modified by LLMs.
+The conceptual arguments, experimental design, analysis, and final conclusions presented in this paper are the original work of the authors.
+We take full responsibility for all content of this work.
+Appendix B Extended Related Work
+Our work focuses on memory and personalization in embodied agents.
+Here, we broaden the related work to include studies on memory-augmented agents, personalization, and memory architectures within the wider research context.
+Memory-augmented agents have increasingly been studied as a means to support long-term reasoning, planning, and adaptive behavior in LLM agents.
+Park et al.
+(2023) propose Generative Agents, which simulate human-like behavior by maintaining a memory stream of past experiences in natural language.
+This memory stream enables agents to reflect on prior events, retrieve relevant information, and plan future actions based on their histories.
+In a similar direction, Xu et al.
+(2025) introduce A-Mem, a dynamic memory system in which memory is represented as evolving and interconnected notes.
+The agent can generate, retrieve, and update these notes over time, thereby supporting adaptability.
+Memory for dialogue and personalization has emerged as a key direction, highlighting how memory supports user-specific context and long-term consistency in interactions.
+Li et al.
+(2025a) present a system that leverages both short-term and long-term memories to maintain user-specific context across multiple conversation sessions, reporting improvements in dialogue coherence and persona consistency.
+Meanwhile, Wang et al.
+(2024a) propose the Self-Controlled Memory (SCM) framework, where the agent dynamically decides when and what to store or retrieve, achieving greater coherence and knowledge retention over extended interactions.
+Alongside these approaches, structured long-term and episodic memory have been introduced to enhance personalization.
+Zhong et al.
+(2023) develop MemoryBank, which stores and retrieves interaction history to support consistent responses across turns.
+Das et al.
+(2024) present Larimar, a framework that enables selective recall and forgetting, providing controllable knowledge updates that can support personalization at scale.
+Architectural perspectives on memory focus not only on building memory components for individual tasks but on treating memory as a fundamental part of the overall system.
+Packer et al.
+(2024) illustrate this idea with MemGPT, which frames memory management through an operating-system analogy.
+In this view, the LLM functions like an operating system that autonomously organizes and manages both its in-context and external memory, rather than relying solely on user-provided prompts.
+This perspective emphasizes memory as a central architectural layer, enabling scalable and autonomous agents that can handle complex, long-horizon tasks.
+Appendix C Details of Evaluation Setup
+C.1 LLM-Powered Embodied Agent Architecture
 Following Szot et al.
-[45], Puig et al.
-[38], Chang et al.
-[7], we adopt a two-layer hierarchical control architecture for our LLM-powered embodied agent.
+(2021); Puig et al.
+(2023); Chang et al.
+(2024), we adopt a two-layer hierarchical control architecture for our LLM-powered embodied agent.
 We utilize the LLM as a high-level policy planner that selects appropriate skills from the predefined skill library.
@@
 Semantic memory.
-For semantic memory, we implement a scene-graph style hierarchical representation, which has demonstrated effectiveness for planning problems [1, 39, 16].
+For semantic memory, we implement a scene-graph style hierarchical representation, which has demonstrated effectiveness for planning problems (Agia et al., 2022; Rana et al., 2023; Gu et al., 2024).
 Following Chang et al.
-[7], we utilize a multi-edge directed graph with three distinct levels to represent environmental entities.
+(2024), we utilize a multi-edge directed graph with three distinct levels to represent environmental entities.
 The top level contains a single root node representing the house environment, the second level comprises room nodes, and the third level encompasses furniture, objects, and agents.
@@
 Episodic memory.
-Our episodic memory is configured to store the ReAct-style formatting that guides the LLMs’ reasoning process [56].
+Our episodic memory is configured to store the ReAct-style formatting that guides the LLMs’ reasoning process (Yao et al., 2023).
 This memory structure captures both the user’s instruction and the complete sequence of <Thought, Action, Observation> triplets generated during task execution.
@@
 Memory retrieval.
-For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer [40] and use the current task instruction as the query.
+For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer (Reimers & Gurevych, 2019) and use the current task instruction as the query.
 To avoid ambiguous or irrelevant memories, we retrieve candidate memories only from the history within the same scene as the current task.
-LLM Setup.
+LLM inference setup.
 We configured the language model with a temperature of 0 to ensure deterministic outputs. For sampling parameters, we set top_p to 1 and top_k to 50.
-B.3 Computing resources
-Our experiments primarily utilized commercial API services rather than local computing resources.
-We used OpenAI’s Chat API for accessing GPT-4o, Claude models through Anthropic’s API, and OpenRouter’s API service for accessing Llama-3.1 [33], Qwen-2.5 [54].
-For running simulation environment we used 8 NVIDIA GeForce RTX 3090 GPUs.
-For our implementation and evaluation, we use Huggingface library2, vLLM library.
-Both libraries are licensed under Apache License, Version 2.0.
-And we used langchain library, under MIT License.
-We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
-Appendix C Details of Memento
-C.1 Personalized Knowledge
+C.2 Personalized Knowledge Details
 We categorize knowledge about personal items as object semantics and knowledge about consistent behaviors as user patterns to structure our evaluation approach.
@@
 Preference A specific way the user prefers to prepare or arrange their environment when a particular situation occurs. my coffee break, cozy decoration spot
-C.2 PartNR Dataset
-PartNR is designed to evaluate planning and reasoning capabilities in embodied tasks and is recognized for its comprehensive collection of natural language instructions in household environments.
-The benchmark includes four primary task types: (1) constraint-free basic rearrangement tasks, (2) spatial tasks requiring reasoning about object positions, (3) temporal tasks with sequential dependencies, and (4) heterogeneous tasks involving actions that can only be performed by human agents.
-We selected PartNR episodes specifically for their complexity beyond simple pick-and-place operations.
-The rich linguistic structure and diverse task requirements make PartNR particularly suitable for evaluating user patterns personalization, enabling us to create scenarios that effectively test an agent’s ability to adapt to user-specific communication patterns.
-C.3 Dataset Statistics
+C.3 Details of Memento
+C.3.1 Foundation Dataset
+We adopt the test set of PartNR (Chang et al., 2024) as our foundation object rearrangement task dataset.
+PartNR is designed to evaluate planning and reasoning capabilities in embodied tasks and includes four primary task types: (1) constraint-free basic rearrangement, (2) spatial tasks requiring reasoning about object positions, (3) temporal tasks with sequential dependencies, and (4) heterogeneous tasks involving human-only actions.
+We select PartNR episodes for their complexity beyond simple pick-and-place operations, filtering out tasks irrelevant to personalized assistance scenarios.
+The rich linguistic structure and diverse task requirements make PartNR particularly suitable for evaluating personalized user patterns, enabling us to create scenarios that effectively test agents’ ability to adapt to user-specific behavioral routines.
+C.3.2 Dataset Generation Process Details
+We leverage GPT-4o and a systematic process to incorporate personalized knowledge into existing task structures, creating tasks tailored for each evaluation stage.
+Captioning process.
+Since we employ LLMs as high-level policy planners for embodied agents, we generate natural language descriptions of objects in the scene using GPT-4o, enabling agents to reason over object properties without relying on direct visual perception.
+We collect object models from the OVMM dataset (Yenamandra et al., 2023) and use GPT-4o to generate descriptions from rendered object images.
+In particular, the Google Scanned dataset (Downs et al., 2022) included within OVMM provides object identities in file names, which we leverage alongside the images to produce more realistic and accurate descriptions.
+Through this process, we generate 1,920 object descriptions spanning 66 object categories.
+The prompts used for generating object descriptions are provided in Appendix F.
+Preprocessing scenes.
+To collect suitable episodes for personalized assistance tasks, we preprocessed and filtered episodes using three criteria.
+We only gathered non-heterogeneous episodes that can be executed by robotic agents, excluding tasks that require human-only actions.
+We filtered out tasks where target objects were not uniquely specified (e.g., "Bring one apple to the kitchen table"), since our task setting requires identification of specific objects based on personalized knowledge that distinguishes them from similar objects.
+When captions for target objects were unavailable, we substituted alternative objects from the same category, or excluded the entire episode if no suitable alternatives existed.
+Distractor sampling.
+To sample distractor objects for episodes focusing on object semantics, we utilized PartNR’s dataset generation methods.
+This approach systematically selects objects located adjacent to target objects on the same receptacles or floor surfaces, requiring agents to differentiate between objects of the same category.
+Figure 10: Episodes with zero success rate (31 in total) were excluded from the analysis.
+Task instruction generation process.
+We prompt GPT-4o to generate personalized knowledge tasks based on the knowledge categories defined in Section C.2.
+For object semantics tasks, we provide target object captions and guide GPT-4o to generate natural semantic descriptions.
+We first instruct the model to create the most plausible subcategory of object semantics, then generate corresponding personalized knowledge.
+The resulting instruction comprises three components: the command instruction, additional object information, and personalized knowledge.
+For example: "Bring the cup on the kitchen table" (command) + "The cup is a white mug with fancy handle" (description) + "That mug is my coffee mug" (personalized knowledge).
+For user pattern tasks, we provide base instructions and allow GPT-4o to infer plausible user patterns corresponding to action sequences.
+Memory acquisition stage instructions concatenate the command with personalized knowledge (e.g., "Bring the cup and dish from the kitchen table to living room.
+That’s my dinner setup").
+Detailed prompts are provided in Appendix F.
+Quality control.
+To ensure dataset quality, we implement a two-stage filtering process after generating episodes for Memento.
+We first apply rule-based heuristic filters guided by six explicit criteria to identify problematic episodes:
+  • Semantic misalignment – Instructions inconsistent with stored memory (e.g., memory states "my favorite cup" but instruction refers to "my cup")
+  • Triviality – Instructions solvable without memory access (e.g., "Prepare my coffee time cup setup on the kitchen table" when setup details are explicitly provided)
+  • Redundancy – Instructions repeating knowledge from previous episodes in the same scene (e.g., "my usual morning cup" when an earlier episode already established "the cup I usually use in the morning")
+  • Environment inconsistency – Instructions conflicting with actual object properties (e.g., describing target cup as red when it’s actually blue, but a distractor cup is red)
+  • Instruction inconsistency – Semantic inconsistency between acquisition and utilization stages (e.g., changing from "pick from A and place at B" to only "place at B")
+  • Temporal irrelevance – Instructions with expired temporal context (e.g., referencing "today’s temperature" when the memory is no longer current)
+Following this rule-based filtering, we test each remaining episode with GPT-4o, excluding episodes where the model fails five consecutive attempts as this indicates potential feasibility or clarity issues.
+This comprehensive quality control process filters out 31 episodes (13.4%) showing consistently poor performance, all with zero success rates indicating complete task failure.
+The remaining 201 episodes demonstrate strong performance with 95% confidence intervals of [0.755, 0.844] for success rate and [0.823, 0.893] for completion rate.
+The high correlation (r = 0.908) between these metrics confirms consistent performance across evaluation measures.
+C.3.3 Inter-Rater Reliability Analysis
+To validate the consistency of our heuristic filtering process, we conduct inter-rater reliability analysis using Krippendorff’s alpha (Krippendorff, 2011).
+Multiple annotators independently evaluate a held-out subset of instructions using our six filtering criteria.
+Since Environment inconsistency and Temporal irrelevance are strictly controlled during instruction generation, they achieve near-perfect agreement and we exclude them from reliability analysis.
+We compute Krippendorff’s alpha for the remaining four categories, obtaining values of 0.75 for Semantic misalignment, 0.87 for Instruction inconsistency, 0.58 for Triviality, and 0.57 for Redundancy.
+The average alpha of 0.69 across these criteria indicates moderate agreement, supporting the reliability of our filtering process for maintaining dataset quality.
+C.3.4 Dataset Statistics
 The dataset comprises 438 episodes, divided into two main stages.
@@
 All episodes were constructed using the Habitat 3.0 simulator.
-C.4 Data Generation Process
-To create stage-specific tasks, we leveraged GPT-4o and systematic process to incorporate personalized knowledge into existing task structures.
-Captioning process.
-Since we employ LLMs as high-level planners for embodied agent, we generated natural language descriptions of the objects at the scene using GPT-4o, to enable agents to reason over object descriptions without relying on direct visual perception.
-We collected object models from the OVMM dataset [57], and used GPT-4o to generate natural language descriptions from rendered object images.
-Especially, the Google Scanned dataset [14], included within OVMM, provides object identities in file names.
-We leveraged this additional information alongside the images to produce more realistic descriptions.
-Through this process, we generated 1,920 object descriptions with 66 categories of objects.
-The prompts used to generate the object descriptions are provided in Appendix E.
-Preprocessing scenes.
-To collect the suitable episodes for our purposes, we preprocessed and filtered episodes.
-First, we only gathered non-heterogeneous episodes, which should be executed by human agents.
-Second, we filtered out tasks where the target object was not uniquely specified (e.g., “Bring one apple to the kitchen table”), as our task setting requires identification and reference of a specific object based on personalized knowledge that distinguishes it from other similar objects.
-Third, in cases where captions for target handles were unavailable, we substituted alternative objects.
-If no objects from the same category were available, we excluded the entire episode from our dataset.
-Distractor sampling.
-To sample distractor objects for episodes focusing on object semantics, we utilized PartNR’s dataset generation methods.
-This approach allowed us to systematically select objects located adjacent to target objects on the same receptacles or floor surfaces, requiring agents to differentiate between objects of the same category.
-Figure 7: Episodes with zero success rate (31 in total) were excluded from the analysis.
-Details of task instruction generation.
-For instruction generation, we prompted GPT-4o to generate personalized knowledge tasks based on the knowledge categories defined in Section C.1.
-For tasks involving object semantics, we provided the captions of the target object pairs along with instructions to guide GPT-4o in generating natural object semantics descriptions.
-(e.g., instruction: "Bring the cup on the kitchen table", object description: "a white mug with fancy handle", personalized knowledge: "The mug is my coffee mug.") We first instructed the model to create the most plausible subcategory of object semantics, which was then used to generate personalized knowledge specific to the objects.
-In these object semantics cases, the instruction comprised three components: command instruction (e.g., "Bring the cup on the kitchen table"), additional information (e.g., "The cup is a white mug with fancy handle."), and personalized knowledge (e.g., "That mug is my coffee mug").
-For user patterns tasks, we provided instructions to GPT-4o and allowed the model to infer the most plausible user patterns corresponding to a sequence of actions.
-For the memory acquisition stage instruction, we also concatenated the command instruction (e.g., "Bring the cup and dish from the kitchen table to living room") with personalized knowledge (e.g., "That’s my dinner setup").
-Detailed information about the prompts we used is provided in Appendix E.
-Quality control.
-After generating episodes for Memento, we implemented a two-stage filtering process for quality control.
-First, we applied heuristic filtering to eliminate episodes with potentially confusing distractor objects in the same scene (e.g., episodes using identical object handles or requiring similar personalized knowledge, which could create ambiguity).
-Second, we tested each episode with GPT-4o and excluded any episode where GPT-4o failed five consecutive attempts, indicating potential issues with task feasibility or clarity.
-(Figure 7) This quality control process resulted in filtering out 31 episodes (13.4% of the total dataset) that exhibited consistently poor performance.
-Specifically, we identified: 31 episodes with zero success rate, indicating complete task failure.
-The remaining 201 episodes showed strong performance metrics, with 95% confidence intervals of [0.755, 0.844] for success rate and [0.823, 0.893] for completion rate.
-The high correlation (r = 0.908) between success rate and completion rate indicates consistent performance across different evaluation metrics.
-These high-quality episodes served as the foundation for generating golden trajectories, which were subsequently used in our memory quality analysis and discussion experiments.
-By establishing GPT-4o’s successful task executions as golden trajectories, we ensure that our evaluation framework is based on demonstrably achievable performance standards, providing a reliable benchmark for assessing memory quality and discussion effectiveness.
-Appendix D Additional Experiments & Analysis
-D.1 Knowledge Type-based Analysis.
-Additional experiment results show the performance of small models—Llama3.1-8b and Qwen2.5-7b—on knowledge type-based analysis (Figure 8).
-Similar to other frontier models, they struggled with utilizing episodic memory on tasks requiring recognition of user patterns.
-For tasks requiring object semantics during the memory acquisition stage, we found that these small models struggle to understand the user’s intent to differentiate objects with DescribeObjectTool and rely primarily on commonsense reasoning, instead of describing the objects.
-Figure 8: Memory acquisition stage and memory utilization stage all models.
-D.2 Knowledge Type-based Analysis on Dual-Memory Tasks
-Figure 9 shows the success rate of each model on dual-memory tasks, compared with the discrete success/failure outcomes when executing each corresponding episode individually during the memory acquisition stage.
-We can observe that almost all models tend to struggle with utilizing user patterns knowledge from different memory sources, with smaller models (Llama3.1-8b and Qwen2.5-7b) demonstrating particularly pronounced difficulties.
-Figure 9: Personalized knowledge type-based analysis on joint-memory tasks, comparing with memory acquisition stage’s corresponding episodes.
-D.3 Memory Quality Analysis
+C.4 Evaluation Method
+We adopt the official evaluation protocol from the PartNR benchmark (Chang et al., 2024), which provides a python-based framework for assessing multi-step rearrangement tasks.
+We use this framework without modification.
+The evaluator analyzes simulator states using three components: (1) propositions that define object relationships to satisfy (e.g., is_on_top([spoon_1], [table_1])), (2) dependencies that define temporal conditions for multi-step instructions (after_satisfied, after_unsatisfied), and (3) constraints that enforce execution requirements (e.g., ordering, object consistency across steps).
+We rely on this system to evaluate tasks involving ambiguous references and sequential dependencies.
+The evaluation produces a percent-complete score and binary success indicator.
+C.5 Computing resources
+Our experiments primarily utilized commercial API services rather than local computing resources.
+We used OpenAI’s Chat API for accessing GPT-4o, Claude models through Anthropic’s API, and OpenRouter’s API service for accessing Llama-3.1 (Meta, 2024), Qwen-2.5 (Yang et al., 2024).
+For running simulation environment we used 8 NVIDIA GeForce RTX 3090 GPUs.
+For our implementation and evaluation, we use Huggingface library2, vLLM library.
+Both libraries are licensed under Apache License, Version 2.0.
+And we used langchain library, under MIT License.
+We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
+Appendix D Experiment & Analysis Details
+In this section, we provide a detailed description of the experimental setup, analysis results, and the memory module, complementing the explanations provided in the main text.
+D.1 User Profile Memory
+Our work introduces a user profile memory module and offers initial evidence that, when combined with episodic memory, structured personalized knowledge can enhance performance on personalized assistance tasks.
+Knowledge graph design.
+We define the user profile memory as a hierarchical knowledge graph that organizes personalized knowledge into three levels—User, Knowledge, and Elements—to provide a structured representation that is both scalable across users and resistant to redundancy from overlapping entities.
+The hierarchical partitioning enables the system to scale to many users while preserving efficiency, since overlapping objects can be represented in a canonical form and connected via edges instead of being redundantly replicated.
+To support this representation, each nodes are connected through two types of edges: hierarchical edges, which capture inclusion relations across and within levels, and temporal edges, which encode the sequential structure of user patterns.
+Together, these edge types allow the graph to comprehensively represent the personalized knowledge that users may require an embodied agent to leverage in real-world interactions.
+Input: User instruction II, Knowledge graph $G$
+Output: Updated knowledge graph $G^{\prime}$
+// Step 1: Extract candidate knowledges from instruction
+$\mathcal{K}\leftarrow ExtractKnowledge(I)$ // $\mathcal{K}=\{(k_{1},t_{1}),(k_{2},t_{2}),\dots\}$
+$C\leftarrow\varnothing$
+1ex
+foreach $(k,t_{k})\in\mathcal{K}$ do
+ // Step 2: Extract elements (patterns/objects/locations) tied to this knowledge
+ $\mathcal{E}\leftarrow ExtractElements(I,k,t_{k})$;
+ // Step 3: Select subgraph by type
+ $G_{t_{k}}\leftarrow Subgraph(G,type=t_{k})$;
+ // Step 4: Retrieve candidates for knowledge and elements
+ $C\leftarrow C\cup SimilaritySearch(G_{t_{k}},k,t_{k})$;
+ foreach $(e,t_{e})\in\mathcal{E}$ do
+    $C\leftarrow C\cup SimilaritySearch(G_{t_{k}},e,t_{e})$;
+// Step 5: Use LLM to decide add or update
+$op\leftarrow LLMResolve(I,C)$;
+if $op=\mathtt{update}$ then
+ return $G^{\prime}\leftarrow UpdateKnowledgeNode(G,Expand(C),I)$;
+else if $op=\mathtt{add}$ then
+ return $G^{\prime}\leftarrow AddKnowledgeNode(G,Expand(C),I)$;
+Algorithm 1 Update Knowledge Graph
+Knowledge graph update procedure.
+While the hierarchical design specifies how knowledge is organized, the graph must also evolve as users provide new information or refine existing entries.
+Algorithm 1 presents the procedure for updating the knowledge graph.
+Given a user instruction $I$ and the current graph $G$, the procedure begins by extracting candidate knowledge $\mathcal{K}$ and associated elements $\mathcal{E}$.
+For each candidate knowledge $k$, the relevant subgraph $G_{t_{k}}$ is filtered, and similarity search is applied to collect matching nodes $C$.
+These candidates are then expanded by collecting all connected edges and their descendant nodes, and the resulting subgraphs are passed to an LLM.
+The LLM determines whether the instruction corresponds to an addition or an update and performs the corresponding operation, yielding the updated graph $G$’.
+Input: User instruction II, User profile memory graph $G$
+Output: Retrieved natural language descriptions $R$
+// Step 1: Parse instruction
+$\mathcal{K}\leftarrow ExtractKnowledge(I)$;
+$\mathcal{C,R}\leftarrow\varnothing$
+// Step 2: Filter subgraph and run similarity search
+foreach $(k,t_{k})\in\mathcal{K}$ do
+ $G_{t}\leftarrow Subgraph(G,type=t)$;
+ $C\leftarrow C\cup SimilaritySearch(G_{t_{k}},k,t_{k})$;
+$C\leftarrow RemoveDuplicates(C)$;
+// Step 3: Reformulate results with LLM
+foreach $(c,t_{c})\in\mathcal{C}$ do
+ $R\leftarrow R\cup Reformulate(Expand(c))$;
+return $R$
+Algorithm 2 RetrievalProcedure
+Retrieval procedure.
+To leverage personalized knowledge during task execution, the agent retrieves information from the user profile memory using a hybrid retrieval pipeline that combines embedding-based search with LLM reasoning.
+This procedure ensures that the retrieved knowledge is both semantically aligned with user instructions and interpretable in natural language.
+Knowledge graph schema. We summarize the node and edge types in Tables 7 and 8. Complete examples of user profile memory instances are shown in Figures 12 and 13.
+Table 7: Edge schema of the user profile memory, including relation types and JSON-style examples.
+Edge Type Source Relation Target Example
+Hierarchical user refers_to knowledge { "source": "u1", "target": "k1", "type": "Hierarchical", "relation": "refers_to" }
+Hierarchical knowledge entails pattern { "source": "k1", "target": "p1", "type": "Hierarchical", "relation": "entails" }
+Hierarchical pattern target object { "source": "p1", "target": "o1", "type": "Hierarchical", "relation": "target" }
+Hierarchical pattern target location { "source": "p1", "target": "l1", "type": "Hierarchical", "relation": "target" }
+Hierarchical knowledge composed_of object { "source": "k1", "target": "o2", "type": "Hierarchical", "relation": "composed_of" }
+Temporal pattern before pattern { "source": "p1", "target": "p2", "type": "Temporal", "relation": "before" }
+Table 8: Node schema of the user profile memory.
+Node Type Attributes Example
+User id, type, name {u1, "user", "james"}
+Knowledge id, type, subtype, alias, description {k_obj1, "knowledge", object_semantics, "my favorite cup", "a yellow cup with a wooden handle"}
+Pattern id, type, name, args {p1, "pattern", "place", [cup, on, livingroom_table]}
+Object id, type, name, granularity {o2, "object", "red dog image cup", instance}
+Location id, type, name, expression {l1, "location", livingroom_table, "on the table"}
+D.2 Success and Error Case Analysis
+As shown in Figure 5, we sampled success and error cases in memory utilization stage.
+Tasks require object semantics knowledge.
+Success cases demonstrate that agents can effectively reference personalized object attributes from episodic memory, correctly identifying and applying the specific information needed for task completion.
+(A.1) However, we observed distinct error patterns when agents needed to utilize this type of knowledge: the most prominent being commonsense reasoning errors (B.1), where agents incorrectly relied on generic object descriptions instead of personalized cues; missed personalization cues (B.2), where agents failed to recognize the need to access personalized knowledge; and memory recall failures (B.3), where agents were unable to locate relevant information despite its presence in the provided context.
+Tasks require user patterns.
+For user patterns tasks, we observed that agents employed two distinct strategies: commonsense reasoning (C.1) treating memory as exemplars for step-by-step reasoning, and direct reference (C.2) for distinctive patterns like "my go-to breakfast." However, both approaches introduced specific vulnerabilities leading to two common failure patterns.
+First, commonsense reasoning (D.1) happened when agents attempted to apply the reasoning-based approach but encountered gaps they couldn’t bridge, leading them to substitute commonsense knowledge that seemed plausible but contradicted established personalized routines.
+Second, inaccurate recall (D.2) occured when agents recognized the need for personalized knowledge but retrieved imprecise or incomplete information from memory.
+The sequential nature of these approaches made them particularly error-prone, as mistakes at any intermediate step propagated through subsequent actions.
+This vulnerability explains the consistently higher failure rates in user patterns tasks across all models, highlighting fundamental challenges in maintaining coherence through multi-step reasoning over episodic memory.
+D.3 Memory Design Experiment Details
+We describe here the experimental setup for Section 6. The following details clarify the episode sampling, memory construction, and retrieval configurations used in this experiment.
+D.3.1 Details of Memory Format Simplification Experiment
+We sampled 30 episodes by selecting 10 episodes from each of three scenes, with careful balancing across task types and difficulty levels informed by our preliminary experimental results.
+To minimize the influence of erroneous trajectories, we provided GPT-4o-generated gold memory as the given episodic memory, thereby reducing the impact of noisy interaction histories.
+We also used GPT-4o to generate summaries of the action trajectories, including user instructions, to facilitate compact memory representations.
+For each evaluation, one exemplar shot was given to the agent, and top-$k$ memory entries were retrieved based on the current query, where $k=10$ for frontier models (GPT-4o and Qwen-2.5-72B) and $k=7$ for smaller models (LLaMA-3.1-8B and Qwen-2.5-7B), considering context length limitations.
+The prompts used for these experiments are provided in Appendix F.3.
+D.3.2 Details of User Profile Memory Experiment
+We conduct experiments by incorporating user profile memory in addition to episodic memory.
+The experiments are conducted with the same models as in Section 6.1 (GPT-4o, Qwen-2.5-72B-Instruct, Llama-3.1-8B-Instruct, and Qwen-2.5-7B-Instruct).
+These models are also employed both to construct the user profile memory and to carry out the retrieving procedure, which combines embedding-based search with LLM reasoning: the LLM parses user instructions and reformulates retrieved graph content into natural language descriptions.
+User profile memory entries are retrieved under a $k=5$ setup without gold memory, while episodic memory is supplied as each agent’s trajectory under a $k=5$ setup with gold memory.
+We evaluate performance on both the single and joint memory utilization stages, and the prompts for memory construction and retrieval are provided in Appendix F.
+D.4 Example Trajectory and Evaluation
+To provide readers with a concrete understanding of our evaluation framework and typical agent performance patterns, we present detailed examples of both successful (Tables 161718) and failure cases (Tables 192021).
+These examples include the complete instruction metadata, agent execution trajectories, and evaluation outcomes, offering insights into when and why agents succeed or struggle with personalized knowledge utilization.
+Appendix E Extended Analysis: Memory Components and Real-World Scenarios
+To provide comprehensive insights into memory utilization for personalized assistance, we conduct extended analysis examining multiple factors affecting agent performance.
+This section covers three primary areas: memory quality effects on agent performance, retrieval strategy effectiveness for episodic memory, and agent behavior in complex real-world scenarios including 3-joint memory tasks and ambiguous instructions.
+E.1 Human Baseline
+Experiment setup.
+To establish a meaningful comparison point for our agent evaluation and validate that our tasks are fundamentally solvable, we conduct a human baseline study.
+We recruit four undergraduate participants to complete a curated set of 20 tasks that represent the challenges our embodied agents face in memory utilization for personalized assistance.
+To ensure comprehensive coverage of task complexity, we categorize task difficulty based on empirical observations of model behavior across our agent evaluations:
+  • Easy: Llama-8B successfully completed the task using its own memory
+  • Medium: Llama-8B failed with its own memory but succeeded when given GPT-4o’s answer memory
+  • Hard: Llama-8B failed even when given GPT-4o’s answer memory
+  • Super hard: GPT-4o failed with its own memory
+  • Random joint-memory tasks
+We implement a text-based interactive interface that mirrors the agent environment while accommodating human interaction patterns.
+Participants issue step-by-step commands and receive observational feedback after each action, similar to how our agents operate.
+Each participant receives a 10-minute tutorial on the interface and completes tasks in the same environment as the agents.
+Crucially, participants have access to the same top-5 retrieved memories that agents receive, allowing direct comparison of memory utilization capabilities.
+Results.
+The human baseline provides both validation and perspective on our evaluation framework.
+All participants successfully completed 100% of the tasks with correct end states, demonstrating that the challenges are fundamentally solvable given appropriate memory access and reasoning capabilities.
+While humans require more time than LLM-based planners, they exhibit superior abilities in referencing episodic memory, recognizing errors, and adapting their actions accordingly.
+This performance establishes a meaningful upper-bound reference for agent capabilities and confirms that our tasks capture genuine challenges in memory utilization for personalized assistance.
+E.2 Memory Quality and Robustness Analysis
+To better understand the factors affecting memory utilization performance, we conduct additional analysis examining two critical aspects of memory-based personalized assistance.
+First, we analyze how memory quality impacts agent performance by comparing gold memory trajectories with retrieved memory from actual agent interactions.
+Second, we evaluate agent robustness when episodic memories are corrupted or incomplete through various noise and degradation conditions.
+These analyses provide insights into the practical limitations and requirements for deploying personalized embodied agents in real-world scenarios where memory quality cannot be guaranteed.
 Figure 11: The results of memory quality analysis.
+E.2.1 Memory Quality Analysis
 Experiment setup.
@@
 Enhancing memory quality through more precise filtering is essential to improve reasoning performance, particularly for complex joint tasks that require integration of multiple memory sources, and especially for smaller models with limited reasoning capabilities.
-D.4 Success and Error Case Analysis
-As shown in Figure 6, we sampled success and error cases in memory utilization stage.
-Tasks require object semantics knowledge.
-Success cases demonstrate that agents can effectively reference personalized object attributes from episodic memory, correctly identifying and applying the specific information needed for task completion.
-(A.1) However, we observed distinct error patterns when agents needed to utilize this type of knowledge: missed personalization cues (B.1), where agents failed to recognize the need to access personalized knowledge; hallucinations (B.2), where agents fabricated non-existent attributes; and memory recall failures (B.3), where agents were unable to locate relevant information despite its presence in the provided context.
-Tasks require user patterns.
-For user patterns tasks, we observed that agents employed two distinct strategies: commonsense reasoning (C.1) treating memory as exemplars for step-by-step reasoning, and direct reference (C.2) for distinctive patterns like "my go-to breakfast." However, both approaches introduced specific vulnerabilities leading to two common failure patterns.
-First, commonsense reasoning (D.1) happened when agents attempted to apply the reasoning-based approach but encountered gaps they couldn’t bridge, leading them to substitute commonsense knowledge that seemed plausible but contradicted established personalized routines.
-Second, inaccurate recall (D.2) occured when agents recognized the need for personalized knowledge but retrieved imprecise or incomplete information from memory.
-The sequential nature of these approaches made them particularly error-prone, as mistakes at any intermediate step propagated through subsequent actions.
-This vulnerability explains the consistently higher failure rates in user patterns tasks across all models, highlighting fundamental challenges in maintaining coherence through multi-step reasoning over episodic memory.
-D.5 Details of Discussion Section Experiments
-Experiment setup details for Section 6.1.
-We sampled 30 episodes by selecting 10 episodes from each of three scenes, with careful balancing across task types and difficulty levels informed by our preliminary experimental results.
-To minimize the influence of erroneous trajectories, we provided GPT-4o-generated gold memory as the given episodic memory, thereby reducing the impact of noisy interaction histories.
-We also used GPT-4o to generate summaries of the action trajectories, including user instructions, to facilitate compact memory representations.
-For each evaluation, one exemplar shot was given to the agent, and top-$k$ memory entries were retrieved based on the current query, where $k=10$ for frontier models (GPT-4o and Qwen-2.5-72B) and $k=7$ for smaller models (LLaMA-3.1-8B and Qwen-2.5-7B), considering context length limitations.
-The prompts used for these experiments are provided in Appendix E.3.
-Experiment setup details for Section 6.2.
-We sampled 30 episodes, selecting 10 episodes from each of three scenes, and tailored them by modifying the personalized knowledge statements and reference components.
-To mitigate potential bias, all episodes were jointly created and validated by two authors.
+E.2.2 Robustness to Noisy and Partial Memories
+Table 9: Model performance across memory acquisition and utilization stages in Memento.
+[…]
+GPT-4o Gold Baseline 14.5 2481.2 85.3 80.6
+Noisy Random 14.6 2759.0 81.5 -3.8 78.6 -2.0
+Shuffle 14.4 2450.9 84.0 -1.3 80.6 0.0
+Partial Random 14.5 2526.9 84.4 -0.9 80.1 -0.5
+w/o Pick 14.6 2555.7 82.9 -2.4 79.1 -1.5
+w/o Place 14.5 2569.8 84.4 -0.9 82.6 +2.0
+w/o Nav 15.4 2526.8 81.4 -3.9 78.5 -2.1
+Experiment setup.
+We conduct additional experiments to evaluate the model’s robustness when episodic memories are corrupted or incomplete.
+As the target model, We leverage GPT-4o, and we organize the experimental setup into two representative degradation types: Noisy Memory and Partial Memory.
+For Noisy Memory, we consider two conditions: (i) random step injection, where unrelated steps from other scenes (30%) are inserted into the memory trajectory to simulate retrieval or perception noise, and (ii) temporal shuffle, where the step order within an episode is randomly permuted to disrupt temporal coherence.
+For Partial Memory, we examine two forms of incompleteness: (i) random step removal, where 20% of steps were randomly removed to test agent robustness under incomplete recall, and (ii) action-type-specific removal, where all steps of a given action type (Place, Place, Navigate) are dropped to analyze their impact on planning.
+Random noise and missing navigation cues degrade performance.
+As shown in Table 9, the model remains robust under most forms of incomplete memory, with only minor degradation.
+However, two factors consistently cause substantial performance drops: (i) the injection of random, noisy entries and (ii) the absence of navigation-related information.
+This indicates that the model is more sensitive to misleading or corrupted inputs than to simple omissions, and that navigation cues play a disproportionately critical role in planning.
+E.3 Retrieval Strategy Analysis
+Effective retrieval of relevant episodic memories is fundamental to successful memory utilization in personalized assistance tasks.
+To understand how retrieval quality affects agent performance, we analyze both baseline retrieval effectiveness and the impact of advanced retrieval strategies on downstream task outcomes.
+We first establish baseline missing rates using similarity-based retrieval to determine optimal retrieval parameters, then examine whether advanced retrieval strategies consistently improve both memory retrieval quality and agent task performance.
+This analysis provides insights into the relationship between retrieval precision and practical task execution in personalized embodied agents.
+Table 10: We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019).
+Retriever Strategy Top-K Missing Rates (%)
+Baseline 3 13.9
+5 3.5
+E.3.1 Baseline Retrieval Analysis.
+To assess whether the corresponding episodic memory is successfully retrieved during memory utilization, we first measure the missing rate under a baseline similarity-based retriever.
+[^1] We adopt an embedding-based method using the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019).
+The results show missing rates of 13.9% for top-3 and 3.5% for top-5, indicating that expanding to top-5 nearly eliminates missing cases.
+Based on this observation, we set top-5 as the default retrieval setup, prioritizing inclusion of the required memory over the risk of distractor entries.
+E.3.2 Advanced Retrieval Strategy Analysis
+Table 11: Performance comparison across different retrieval strategies.
+Model Strategy Top-K Missing Rates (%) Percent Complete $\uparrow$ Success Rate $\uparrow$
+GPT-4o Baseline 3 13.9 82.9 79.5
+5 3.5 84.7 82.1
+Reranker 3 2.5 87.2 84.1
+5 1.5 84.3 80.6
+Query Expansion 3 9.0 84.9 81.1
+5 1.5 86.2 81.1
+Qwen-2.5-72b Baseline 3 13.9 75.3 70.1
+5 3.5 75.4 70.0
+Reranker 3 7.6 73.0 67.2
+5 2.1 70.9 64.1
+Query Expansion 3 8.5 75.6 69.2
+5 4.0 72.6 66.2
+Experiment setup.
+Building on this baseline, we further examine the effect of advanced retrieval strategies on the instruction following capability.
+We compared three approaches: (i) the baseline similarity-based retreival, (ii) reranker-based retrieval, which applies an LLM-based pointwise reranker (Sachan et al., 2022; Sun et al., 2023) to an initial pool of 10 candidates before selecting the final top-3 or top-5, and (iii) query expansion-based retrieval, where queries are reformulated before retrieval (Jagerman et al., 2023; Wang et al., 2023b).
+These strategies are evaluated with both GPT-4o and Qwen-2.5-72b, using the same models to compute reranker scores and produce expanded queries.
+Improved retrieval does not ensure improved outcomes.
+As shown in Table 11, retrieval strategies affect not only memory inclusion but also downstream task performance.
+First, applying reranker or query expansion substantially reduce the missing rate, particularly under the top-3 setting, where improvements range from 4.9pp to 11.4pp.
+This confirms their effectiveness in narrowing the gap between top-3 and top-5 retrieval.
+Second, performance gains arere not consistent across models.
+For example, with Qwen, query expansion under the top-5 setting even degraded retrieval, suggesting that the effectiveness of query expansion critically depends on the quality of the model generating the reformulated queries.
+Finally, except for GPT-4o at top-3, introducing advanced retrieval strategies further reduces downstream performance, likely due to the inclusion of semantically similar but task-irrelevant memories.
+These results suggest that simply increasing retrieval relevance does not necessarily yield better task performance, particularly in real-world contexts with dense and noisy memory pools.
+E.4 Real-World Scenario Analysis
+While our main evaluation demonstrates fundamental limitations in memory utilization for personalized assistance, real-world deployment introduces additional complexities that extend beyond our controlled experimental settings.
+To explore these practical challenges, we conduct three sets of experiments that examine agent behavior under more realistic conditions.
+First, we evaluate scalability by testing agents on 3-joint memory tasks that require integrating knowledge from multiple episodes.
+Second, we assess how agents handle ambiguous instructions that indirectly reference personalized knowledge through contextual cues rather than explicit mentions.
+Finally, we investigate agents’ ability to generalize learned user preferences across different but related contexts, such as applying spatial arrangements from one room to another.
+These experiments provide insights into additional hurdles for successful real-world deployment of personalized embodied agents.
+E.4.1 Real-World Multi-Episode Evaluation
+Table 12: Performance results on multiple joint-memory tasks.
+Model Task PC (%) SR (%)
+GPT-4o joint memory task 86.7 63.9
+GPT-4o 3-joint memory task 61.2 22.0
+Qwen-2.5-72b joint memory task 68.9 36.1
+Qwen-2.5-72b 3-joint memory task 42.4 7.3
+In practice, user requirements are rarely limited to information from just one or two episodes, but often require knowledge distributed across multiple episodes.
+To explore this realistic scenario, we extend our joint-memory task to cases involving three memories, which require agents to effectively integrate three independent memory sources to solve the task.
+Following the same methodology as our joint-memory task construction, we sequentially concatenated three personalized instructions to generate a total of 41 tasks.
+We observed substantial performance drops compared to the two-memory joint-memory tasks for both GPT-4o and Qwen-2.5-72b-instruct.
+This suggests that multi-memory reasoning introduces significant difficulty due to temporal gaps and abstraction challenges.
+We believe this degradation reflects the true complexity of real-world personalized tasks, and addressing this challenge remains critical for future research.
+E.4.2 Agent Behavior under Ambiguous Instructions from Users
+Table 13: Performance under ambiguous queries for personalized knowledge.
+PC (Percent Complete) and SR (Success Rate) (%) indicate how well agents resolve indirect references to personalized knowledge from memory.
+Model PC (%) SR (%)
+GPT-4o (Baseline) 92.0 90.0
+GPT-4o 80.4 73.3
+Qwen-2.5-72b (Baseline) 75.1 66.7
+Qwen-2.5-72b 59.6 53.3
+Experiment setup.
+While Memento focuses on evaluating personalized knowledge grounding with explicit references to prior interactions, real-world human-agent communication often involves ambiguous or indirect references.
+To explore this challenge, we conducted a proof-of-concept experiment to assess whether current models can interpret ambiguous instructions that indirectly refer to previously encountered personalized knowledge.
+We created a tailored set of 30 tasks by sampling 10 episodes from each of three scenes and modifying them to include indirect references to personalized knowledge through contextual cues, synonyms, or causal references (e.g., Can you set my afternoon tea time routine?
+$\rightarrow$ I’m about to enjoy my afternoon tea.
+Could you set things up as I like them?).
+To ensure quality and mitigate potential bias, all episodes were jointly created and validated by two authors.
 For each model, we first generated memory traces during the memory acquisition stage and then evaluated memory usage during the utilization stage.
@@
 In each case, the user rephrases the original request to refer indirectly to previously established personalized knowledge, simulating natural variability in real-world human-agent communication.
+Result.
+The results, shown in Table 13, reveal the degradation in performance, indicating that handling ambiguous queries remains a key challenge for future personalized embodied agents.
+We view this as a promising direction for future work, where we plan to systematically extend Memento to evaluate LLM-powered embodied agents’ capabilities under ambiguous and implicit reference scenarios, aiming to better reflect the complexity of real-world human-agent interaction.
 Table 14: Examples of ambiguous instruction modifications in the additional experiment dataset.
@@
 Could you help me tidy up by moving my book received to celebrate completing my studies, the candle with a brown and blue gradient, and the round white clock with black numbers to the stand?
-Appendix E Prompts
-E.1 Prompts for Dataset Generation
-E.2 Prompts for Agent
-E.3 Prompts for Discussion
-Appendix F Extended Related Work
-Recent research has increasingly emphasized the integration of memory mechanisms into large language model (LLM) agents to support long-term reasoning, planning, and personalization.
-Park et al.
-[36] propose Generative Agents, which simulate human-like behavior by maintaining a memory stream of past experiences in natural language.
-This enables agents to reflect, retrieve, and plan based on their individual histories.
-Similarly, Xu et al.
-[53] introduce A-Mem, a dynamic memory system inspired by Zettelkasten, which structures memory as evolving and interconnected notes that the agent can generate, retrieve, and update over time, supporting agentic autonomy and adaptability.
-Personalization in dialogue agents has also been explored through memory-enhanced architectures.
-Li et al.
-[27] present a personalized dialogue agent that leverages both short-term and long-term memories to maintain user-specific context across sessions, significantly improving response consistency and contextual relevance.
-Wang et al.
-[46] propose the Self-Controlled Memory (SCM) framework, where an agent dynamically decides when and what to store or retrieve from memory, leading to improved coherence and knowledge retention over extended interactions.
-Zhong et al.
-[60] introduce MemoryBank, a structured long-term memory module that enhances LLMs by storing and retrieving relevant interaction history to support consistent and personalized user responses across multiple turns.
-Das et al.
-[12] present Larimar, a framework that integrates episodic memory control into LLMs, enabling selective recall and forgetting to improve memory scalability and privacy-aware reasoning.
-More broadly, Packer et al.
-[35] conceptualize memory management as an OS-level abstraction with MemGPT, enabling an LLM to autonomously manage and interact with its internal and external memory hierarchies.
-This work highlights how memory can act as a core architectural layer to enable scalable, autonomous agents capable of long-horizon tasks and continuous learning.
-Appendix G License
-For our implementation and evaluation, we use Huggingface library [^4] https://huggingface.co/ and vLLM library.
-Both libraries are licensed under Apache License, Version 2.0.
-We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
-G.1 License for the Assets
-The existing assets used in this research are properly credited and their licenses respected:
-G.2 License for the Codes
-The existing codes used in this research are properly credited and their licenses respected:
+E.4.3 Personalized Knowledge Generalization
+Experiment setup.
+A critical aspect of personalized embodied agent is their ability to generalize learned user preferences across different but related contexts.
+This includes applying knowledge about a user’s breakfast routine to dinner preparation tasks, or transferring spatial preferences from one room to another.
+To investigate this capability, we designed generalization tasks that test whether embodied agents can abstract and reuse personalized knowledge as transferable templates.
+We extended our user pattern tasks by sampling existing scenarios and systematically altering target locations, thereby evaluating whether agents can adapt user preferences to novel but structurally similar situations.
+To encourage explicit memory utilization, we incorporated contextual cues in the instructions, such as "on [another receptacle] this time," clearly signaling that established preferences should apply in the modified context.
+Our generalization tasks were constructed using relatively simple base cases that GPT-4o consistently solves in standard settings.
+Table 15: Agent performance on personalized knowledge generalization tasks.
+Model Task PC (%) SR (%)
+GPT-4o Baseline 100.0 100.0
+GPT-4o Generalization 92.5 83.3
+Qwen-2.5-72b Baseline 75.0 75.0
+Qwen-2.5-72b Generalization 85.8 75.0
+Result.
+As in Table 15, we observed that GPT-4o occasionally fails by defaulting to commonsense reasoning rather than retrieving and adapting the relevant personalized memory.
+Interestingly, Qwen-2.5-72b demonstrated improved completion rates on these tasks.
+We attribute this improvement to the explicit instruction cues, which likely reduced ambiguity and guided the model toward memory-based reasoning.
+While these preliminary experiments provide initial insights into generalization capabilities, they represent an important step toward understanding how agents can flexibly apply personalized knowledge.
+Future work could build on these initial observations to better understand how agents transfer personalized knowledge across different contexts.
+Appendix F Prompts
+F.1 Prompts for Dataset Generation
+F.2 Prompts for Agent
+F.3 Prompts for Memory Design Experiment
+Table 16: Episode metadata for the example task.
+Field Value
+episode_type user_pattern
+subtype preference
+task Could you decorate the bedroom as I like?
+task_percent_complete 1.0
+task_state_success 1.0
+Table 17: Trajectory for the example task.
+# Thought Action Result / State $\Delta$
+1 Since there are no objects found yet, I should start by exploring the bedroom to locate any objects that can be used for decoration. Explore
+(bedroom_1)
+No objects found
+2 Since no objects were found in bedroom_1, I should explore another room to locate objects that can be used for decoration.
+I will explore the living room next, as it is likely to contain decorative items.
+Navigate
+(to living_room_1)
+Location changed
+3 I have found a plant container and a statue on the shelves in the living room, which can be used for decoration in the bedroom. I will start by picking up the plant container from the shelves. Pick
+(plant_container_1)
+plant_container_1 held
+4 Now that I am close to the shelves, I will attempt to pick up the plant container to move it to the bedroom for decoration. (continue) Success
+5 I have successfully picked up the plant container. Now, I need to navigate to the bedroom to place it for decoration. Navigate
+(to bedroom_1)
+Arrived in bedroom_1
+6 I have successfully navigated to bedroom_1. Now, I need to find a suitable location to place the plant container for decoration. I will place it on the chest of drawers in the bedroom. Place
+(plant_container_1
+on chest_of_drawers_54)
+Object placed
+7 Now that I am close to chest_of_drawers_54, I will place the plant container on it for decoration. (continue) Confirmed placement
+8 I have successfully placed the plant container on chest_of_drawers_54 in bedroom_1.
+Now, I need to return to the living room to pick up the statue from the shelves for further decoration in the bedroom.
+Navigate
+(to living_room_1)
+Location changed
+9 I am back in the living room. I need to navigate closer to the shelves to pick up the statue for further decoration in the bedroom. Positioning for pick
+10 Now that I am close to the shelves, I will attempt to pick up the statue to move it to the bedroom for decoration. Pick
+(statue_0)
+statue_0 held
+11 I have successfully picked up the statue. Now, I need to navigate to the bedroom to place it for decoration. Navigate
+(to bedroom_1)
+Location changed
+12 I have successfully navigated to bedroom_1.
+Now, I need to find a suitable location to place the statue for decoration.
+I will place it on the chest of drawers next to the plant container.
+Positioning
+13 Now that I am close to chest_of_drawers_54, I will place the statue next to the plant container for decoration. Place
+(statue_0 on
+chest_of_drawers_54
+next_to plant_container_1)
+Object placed
+Table 18: Evaluation functions for the example task.
+# Function Args Description
+1 is_on_top object_handles=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], receptacle_handles=[bedroom_chest_of_drawers_handles],
+number=1
+Plant container is on drawer
+2 is_on_top object_handles=[Dog_Statue_4_:0000],
+receptacle_handles=[bedroom_chest_of_drawers_handles],
+number=1
+Statue is on drawer
+3 is_next_to entity_handles_a=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], entity_handles_b=[Dog_Statue_4_:0000],
+number=1,
+l2_threshold=1.0
+Plant and statue are next to each other
+Table 19: Episode metadata for the example task.
+Field Value
+episode_type user_pattern
+subtype routine
+task Could you set up the candle for a cozy evening?
+task_percent_complete 0
+task_state_success 0
+Table 20: Trajectory for the example task.
+# Thought Action Result / State $\Delta$
+1 Since there are no objects found yet, I should start by exploring the tv room to locate a candle that can be set up for a cozy evening. Navigate
+(to tv_1)
+Location changed
+2 I have found a candle, candle_0, on table_34 in the tv room. To set it up for a cozy evening, I will first need to pick it up from the table. Pick
+(candle_0)
+candle_0 held
+3 Now that I am close to table_34, I will attempt to pick up the candle, candle_0, to set it up for a cozy evening. (continue) Success
+4 I have successfully picked up the candle, candle_0. To set it up for a cozy evening, I will place it on a suitable surface in the tv room, such as the tv table, table_14. Place
+(candle_0 on table_14)
+Object placed
+5 I am now close to table_14 in the tv room. I will place the candle, candle_0, on this table to set it up for a cozy evening. (continue) Failed to meet task constraints
+Table 21: Evaluation functions for the example task.
+# Function Args Description
+1 is_on_top object_handles=[candle], receptacle_handles=[chair_handles],
+number=1
+Candle must be on chair
+2 is_on_top object_handles=[candle], receptacle_handles=[table_handles],
+number=1
+Candle must be on table (after chair)
+3 Constraint TemporalConstraint: Placing the candle on the chair before placing it on the table.
+Figure 12: Example of a user profile memory graph representing object semantics knowledge.
+Figure 13: Example of a user profile memory graph representing user pattern knowledge.

### v3

@@
 Given our observation that episodic memory provides both personalized knowledge and in-context learning benefits, we design a hierarchical knowledge graph-based user-profile memory module that separately manages personalized knowledge, achieving substantial improvements on both single and joint-memory tasks.
-Project website: https://connoriginal.github.io/MEMENTO
+Our code and data is available at https://github.com/Connoriginal/MEMENTO.
 1 Introduction
@@
 Personalization for embodied agents.
-The importance of personalization in robotics has long been recognized (Dautenhahn, 2004; Lee et al., 2012; Clabaugh & Matarić, 2018), particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users.
-Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement (Kapelyukh & Johns, 2021; Wu et al., 2023), table settings (Puig et al., 2021), or personalized object navigations (Dai et al., 2024; Barsellotti et al., 2024).
+The importance of personalization in robotics has long been recognized (Dautenhahn, 2004; Lee et al., 2012; Clabaugh and Matarić, 2018), particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users.
+Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement (Kapelyukh and Johns, 2021; Wu et al., 2023), table settings (Puig et al., 2021), or personalized object navigations (Dai et al., 2024; Barsellotti et al., 2024).
 Recently, Xu et al.
@@
 We formulate the personalized object rearrangement task as a Partially Observable Markov Decision Process (POMDP) $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor.
-At timestep $t$, the agent receives a text observation $w_{t}\in\Omega$ describing nearby objects and generates actions via LLM policy $\pi$.
+At timestep $t$, the agent receives a text observation $w_{t}\in\Omega$ describing nearby objects based on the previous action $a_{t-1}$ and generates action $a_{t}$ via LLM policy $\pi$.
 Given a natural language instruction $I$ that may contain personalized references (e.g., "Place my favorite mug on the table"), the agent must derive the goal representation $g$ using both the instruction and personalized knowledge from memory: $\phi(I,M)\longrightarrow g={(o_{i},l_{i})}_{i=1}^{k}$, where $M={h_{1},h_{2},\ldots,h_{n}}$ represents the set of episodic memories from previous interactions, and each $(o_{i},l_{i})$ pair specifies target object $o_{i}$ and location $l_{i}$.
@@
 Retrieval setup.
-For memory retrieval in the utilization stage, we use top-$k$ similarity-based retrieval with all-mpnet-base-v2 (Reimers & Gurevych, 2019), treating the current instruction as a query and episodic memory instructions as keys.
+We first run the memory acquisition stage for all episodes to accumulate episodic memories, which are then used in the memory utilization stage.
+For memory retrieval in the utilization stage, we use top-$k$ similarity-based retrieval with all-mpnet-base-v2 (Reimers and Gurevych, 2019), treating the current instruction as a query and episodic memory instructions as keys.
 Since our primary objective is evaluating memory utilization capability rather than retrieval performance, we ensure the corresponding gold memories are always included in the top-$k$ results through randomly replacing retrieved candidates when necessary.
 We set $k=5$ as our default configuration because this achieves 96.5% recall for gold memories in realistic retrieval scenarios, compared to only 86.1% recall at $k=3$.
+Comprehensive analysis of various retrieval strategies without gold memory guarantees is provided in Appendix E.3.
 4 Evaluating Memory Utilization Across Embodied Agents
@@
 We found that agents increasingly rely on commonsense knowledge rather than personalized memories as retrieved memories increase, especially on user patterns, suggesting that embodied agents suffer from information overload.
-Therefore, even though LLMs demonstrate impressive capabilities in processing long-context information (Wang et al., 2024b; Liu et al., 2025), memory systems that retrieve appropriate numbers of memories are crucial for personalized embodied agents.
-We provide further analysis of retrieval strategies in scenarios without guaranteed gold memory access in Appendix E.3.
+Therefore, even though LLMs demonstrate impressive capabilities in processing long-context information (Wang et al., 2024b; Liu et al., 2025), which we confirm in Appendix D.5, memory systems that retrieve appropriate numbers of memories are crucial for personalized embodied agents.
 5.2 Investigating Agent Behavior in Joint-memory Tasks
@@
 We hope these findings will inspire future research toward more sophisticated memory architectures that enable truly effective personalized embodied agents.
+Acknowledgements
+This project is supported by Microsoft Research Asia.
+This research was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT), under the Global Research Support Program in the Digital Field program (RS-2024-00436680), National AI Research Lab Project (RS-2024-00457882), Artificial Intelligence Graduate School Program (Yonsei University) (RS-2020-II201361).
+We also thank Donghyun Kim, Juhyun Park, Sungyong Park, and Hyeonjong Ju for their helpful discussions and support.
+Jinyoung Yeo is the corresponding author.
 Ethics Statement
@@
 To mitigate these issues, future systems should adopt on-device processing, minimize retention, require user consent, and apply secure, transparent memory handling.
-We will release only de-identified or synthetic assets and provide documentation of safeguards to reduce misuse risk.
 Reproducibility Statement
 To facilitate full reproducibility of our results, we include complete evaluation dataset details and experimental procedures in the appendices.
-Furthermore, we publish our code at https://anonymous.4open.science/r/MEMENTO with thorough documentation.
+Furthermore, we publish our code at  https://github.com/Connoriginal/MEMENTO with thorough documentation.
 Appendix A Limitations, Privacy, Societal Impacts and LLM Usage
@@
 This capability has the potential to enhance convenience, efficiency, and user satisfaction in everyday environments—benefiting a wide range of users regardless of age or ability.
-However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents (Rueben & Smart, 2016; Lee et al., 2012).
+However, personalization also raises concerns about privacy, bias reinforcement, and over-dependence on AI agents (Rueben and Smart, 2016; Lee et al., 2012).
 Since episodic memory involves storing interaction history, future systems must consider secure and transparent memory handling.
@@
 Memory retrieval.
-For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer (Reimers & Gurevych, 2019) and use the current task instruction as the query.
+For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer (Reimers and Gurevych, 2019) and use the current task instruction as the query.
 To avoid ambiguous or irrelevant memories, we retrieve candidate memories only from the history within the same scene as the current task.
@@
 C.3.3 Inter-Rater Reliability Analysis
+Figure 11: Distribution of required (object, location) pairs across object semantics and user pattern tasks.
 To validate the consistency of our heuristic filtering process, we conduct inter-rater reliability analysis using Krippendorff’s alpha (Krippendorff, 2011).
@@
 All episodes were constructed using the Habitat 3.0 simulator.
+Number of (target, location) pairs.
+We compared the number of (object, location) pairs required for successful completion across both object semantics and user pattern tasks.
+As shown in Figure 11, the distribution of required (object, location) pairs is nearly identical between the two personalized knowledge types.
+This indicates that the performance difference between object semantics and user patterns stems from the intrinsic characteristics of each knowledge type rather than differences in task complexity measured by the number of required object-location pairs.
 C.4 Evaluation Method
@@
 Together, these edge types allow the graph to comprehensively represent the personalized knowledge that users may require an embodied agent to leverage in real-world interactions.
+Knowledge graph update procedure.
+While the hierarchical design specifies how knowledge is organized, the graph must also evolve as users provide new information or refine existing entries.
+Algorithm 1 presents the procedure for updating the knowledge graph.
+Given a user instruction $I$ and the current graph $G$, the LLM processes the instruction by extracting candidate knowledge $\mathcal{K}$ and associated elements $\mathcal{E}$.
+For each candidate knowledge $k$, the relevant subgraph $G_{t_{k}}$ is filtered, and similarity search is applied to collect matching nodes $C$.
+These candidates are then expanded by collecting all connected edges and their descendant nodes, and the resulting subgraphs are passed to an LLM.
+The LLM determines whether the instruction corresponds to an addition or an update, executes the corresponding operation.
+For an update, the LLM removes the existing knowledge node and generates a new one, and for an addition it simply creates a new knowledge node.
+In both cases, the LLM determines whether to reuse or instantiate the associated object, location nodes based on whether their similarity scores exceed a predefined threshold.
 Input: User instruction II, Knowledge graph $G$
@@
     $C\leftarrow C\cup SimilaritySearch(G_{t_{k}},e,t_{e})$;
-// Step 5: Use LLM to decide add or update
-$op\leftarrow LLMResolve(I,C)$;
+// Step 5: Decide add or update
+$op\leftarrow DecideAddOrUpdate(I,C)$;
 if $op=\mathtt{update}$ then
@@
 Algorithm 1 Update Knowledge Graph
-Knowledge graph update procedure.
-While the hierarchical design specifies how knowledge is organized, the graph must also evolve as users provide new information or refine existing entries.
-Algorithm 1 presents the procedure for updating the knowledge graph.
-Given a user instruction $I$ and the current graph $G$, the procedure begins by extracting candidate knowledge $\mathcal{K}$ and associated elements $\mathcal{E}$.
-For each candidate knowledge $k$, the relevant subgraph $G_{t_{k}}$ is filtered, and similarity search is applied to collect matching nodes $C$.
-These candidates are then expanded by collecting all connected edges and their descendant nodes, and the resulting subgraphs are passed to an LLM.
-The LLM determines whether the instruction corresponds to an addition or an update and performs the corresponding operation, yielding the updated graph $G$’.
 Input: User instruction II, User profile memory graph $G$
@@
 This procedure ensures that the retrieved knowledge is both semantically aligned with user instructions and interpretable in natural language.
-Knowledge graph schema. We summarize the node and edge types in Tables 7 and 8. Complete examples of user profile memory instances are shown in Figures 12 and 13.
+Knowledge graph schema. We summarize the node and edge types in Tables 7 and 8 and provide complete examples of user profile memory instances in Figures 12 and 13.
 Table 7: Edge schema of the user profile memory, including relation types and JSON-style examples.
@@
 Location id, type, name, expression {l1, "location", livingroom_table, "on the table"}
+Figure 12: Example of a user profile memory graph representing object semantics knowledge.
+Figure 13: Example of a user profile memory graph representing user pattern knowledge.
 D.2 Success and Error Case Analysis
@@
 These examples include the complete instruction metadata, agent execution trajectories, and evaluation outcomes, offering insights into when and why agents succeed or struggle with personalized knowledge utilization.
+Table 9: Episode metadata for the example task.
+Field Value
+episode_type user_pattern
+subtype preference
+task Could you decorate the bedroom as I like?
+task_percent_complete 1.0
+task_state_success 1.0
+Table 10: Trajectory for the example task.
+# Thought Action Result / State $\Delta$
+1 Since there are no objects found yet, I should start by exploring the bedroom to locate any objects that can be used for decoration. Explore
+(bedroom_1)
+No objects found
+2 Since no objects were found in bedroom_1, I should explore another room to locate objects that can be used for decoration.
+I will explore the living room next, as it is likely to contain decorative items.
+Navigate
+(to living_room_1)
+Location changed
+3 I have found a plant container and a statue on the shelves in the living room, which can be used for decoration in the bedroom. I will start by picking up the plant container from the shelves. Pick
+(plant_container_1)
+plant_container_1 held
+4 Now that I am close to the shelves, I will attempt to pick up the plant container to move it to the bedroom for decoration. (continue) Success
+5 I have successfully picked up the plant container. Now, I need to navigate to the bedroom to place it for decoration. Navigate
+(to bedroom_1)
+Arrived in bedroom_1
+6 I have successfully navigated to bedroom_1. Now, I need to find a suitable location to place the plant container for decoration. I will place it on the chest of drawers in the bedroom. Place
+(plant_container_1
+on chest_of_drawers_54)
+Object placed
+7 Now that I am close to chest_of_drawers_54, I will place the plant container on it for decoration. (continue) Confirmed placement
+8 I have successfully placed the plant container on chest_of_drawers_54 in bedroom_1.
+Now, I need to return to the living room to pick up the statue from the shelves for further decoration in the bedroom.
+Navigate
+(to living_room_1)
+Location changed
+9 I am back in the living room. I need to navigate closer to the shelves to pick up the statue for further decoration in the bedroom. Positioning for pick
+10 Now that I am close to the shelves, I will attempt to pick up the statue to move it to the bedroom for decoration. Pick
+(statue_0)
+statue_0 held
+11 I have successfully picked up the statue. Now, I need to navigate to the bedroom to place it for decoration. Navigate
+(to bedroom_1)
+Location changed
+12 I have successfully navigated to bedroom_1.
+Now, I need to find a suitable location to place the statue for decoration.
+I will place it on the chest of drawers next to the plant container.
+Positioning
+13 Now that I am close to chest_of_drawers_54, I will place the statue next to the plant container for decoration. Place
+(statue_0 on
+chest_of_drawers_54
+next_to plant_container_1)
+Object placed
+Table 11: Evaluation functions for the example task.
+# Function Args Description
+1 is_on_top object_handles=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], receptacle_handles=[bedroom_chest_of_drawers_handles],
+number=1
+Plant container is on drawer
+2 is_on_top object_handles=[Dog_Statue_4_:0000],
+receptacle_handles=[bedroom_chest_of_drawers_handles],
+number=1
+Statue is on drawer
+3 is_next_to entity_handles_a=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], entity_handles_b=[Dog_Statue_4_:0000],
+number=1,
+l2_threshold=1.0
+Plant and statue are next to each other
+Table 12: Episode metadata for the example task.
+Field Value
+episode_type user_pattern
+subtype routine
+task Could you set up the candle for a cozy evening?
+task_percent_complete 0
+task_state_success 0
+Table 13: Trajectory for the example task.
+# Thought Action Result / State $\Delta$
+1 Since there are no objects found yet, I should start by exploring the tv room to locate a candle that can be set up for a cozy evening. Navigate
+(to tv_1)
+Location changed
+2 I have found a candle, candle_0, on table_34 in the tv room. To set it up for a cozy evening, I will first need to pick it up from the table. Pick
+(candle_0)
+candle_0 held
+3 Now that I am close to table_34, I will attempt to pick up the candle, candle_0, to set it up for a cozy evening. (continue) Success
+4 I have successfully picked up the candle, candle_0. To set it up for a cozy evening, I will place it on a suitable surface in the tv room, such as the tv table, table_14. Place
+(candle_0 on table_14)
+Object placed
+5 I am now close to table_14 in the tv room. I will place the candle, candle_0, on this table to set it up for a cozy evening. (continue) Failed to meet task constraints
+Table 14: Evaluation functions for the example task.
+# Function Args Description
+1 is_on_top object_handles=[candle], receptacle_handles=[chair_handles],
+number=1
+Candle must be on chair
+2 is_on_top object_handles=[candle], receptacle_handles=[table_handles],
+number=1
+Candle must be on table (after chair)
+3 Constraint TemporalConstraint: Placing the candle on the chair before placing it on the table.
+D.5 LLMs Long Context Comprehension Analysis
+To investigate whether the information overload bottleneck in LLM-powered embodied agents for personalized assistance stems from limitations in processing long-context information, we conduct an experiment in a QA format to evaluate whether LLMs can properly reference personalized information from long contexts.
+Experiment setup.
+We provide LLMs with the episodic memory information required for personalized assistance tasks along with the corresponding task instructions.
+We then prompted the models: "Identify the personalized knowledge necessary to perform this instruction based on the context provided below." To evaluate the responses, we use GPT-4o as an LLM-as-a-Judge, supplying it with the stored gold memory and task instruction as references to assess whether the LLM’s identified personalized knowledge aligns with that in the gold memory.
+Table 15: Accuracy on long-context QA with varying numbers of retrieved episodic memories (Top-$k$).
+Model k = 1 3 5 7 10 20
+GPT-4o 100% 98.71% 98.28% 96.98% 96.98% 95.69%
+Qwen-2.5-72b 100% 93.97% 94.40% 93.53% 93.10% 83.19%
+Results.
+As shown in the Table 15, all models demonstrate high accuracy in correctly identifying personalized knowledge in the QA format.
+Compared to the results in Section 5.1 Figure 6, this indicates that current LLMs have the ability to reference personalized information from long contexts, but encounter difficulties when applying this personalized information to planning tasks.
 Appendix E Extended Analysis: Memory Components and Real-World Scenarios
@@
 To better understand the factors affecting memory utilization performance, we conduct additional analysis examining two critical aspects of memory-based personalized assistance.
-First, we analyze how memory quality impacts agent performance by comparing gold memory trajectories with retrieved memory from actual agent interactions.
+First, we analyze how memory quality impacts agent performance by comparing high quality memory trajectories with acquired memory trajectories from actual agent interactions.
 Second, we evaluate agent robustness when episodic memories are corrupted or incomplete through various noise and degradation conditions.
 These analyses provide insights into the practical limitations and requirements for deploying personalized embodied agents in real-world scenarios where memory quality cannot be guaranteed.
-Figure 11: The results of memory quality analysis.
+Table 16: Performance comparison of models when using high-quality memory versus acquired (self-acquired) memory, evaluated in both single- and joint-memory settings.
+Model Memory Stage Replanning Count $\downarrow$ Sim Steps $\downarrow$ PC (%) $\uparrow$ SR (%) $\uparrow$
+GPT-4o High-quality Single Memory 14.5 2481.2 85.3 80.6
+Joint Memory 25.8 3619.4 83.3 63.9
+Baseline Single Memory 16.1 2450.8 87.9 85.1
+Joint Memory 28.9 3480.7 86.7 63.9
+Claude-3.5-Sonnet High-quality Single Memory 14.1 2266.4 77.2 70.1
+Joint Memory 24.2 3226.6 69.9 38.9
+Baseline Single Memory 15.3 2258.8 69.3 63.7
+Joint Memory 27.8 3198.8 64.6 33.3
+Qwen-2.5-72B High-quality Single Memory 15.2 2630.4 74.7 68.2
+Joint Memory 26.4 4082.1 64.8 38.9
+Baseline Single Memory 17.5 2691.2 72.6 67.2
+Joint Memory 31.3 4027.1 68.9 36.1
+Llama-70B High-quality Single Memory 15.1 2364.0 72.6 67.2
+Joint Memory 27.1 3374.5 64.1 27.8
+Baseline Single Memory 19.0 2566.6 72.2 66.7
+Joint Memory 31.4 3425.2 51.3 8.3
+Llama-8B High-quality Single Memory 16.9 2739.0 65.1 59.2
+Joint Memory 26.4 3753.2 54.0 13.9
+Baseline Single Memory 19.0 3131.7 48.1 35.0
+Joint Memory 27.4 3478.2 35.3 8.3
+Qwen-2.5-7B High-quality Single Memory 15.7 2991.5 55.7 46.3
+Joint Memory 23.9 4402.0 34.0 8.3
+Baseline Single Memory 21.8 3271.4 39.1 27.4
+Joint Memory 26.9 4148.6 33.7 5.6
 E.2.1 Memory Quality Analysis
 Experiment setup.
-We further analyze the effect of memory quality by comparing gold memory, consisting of successful and shortest-path trajectories that serves as high-quality references with the memory obtained from interaction histories in the declaration stage.
+We further analyze the effect of memory quality by comparing high quality memory, consisting of successful and shortest-path trajectories that serves as curated references with the memory obtained from interaction histories in the memory acquisition stage, with predicted memory obtained via automated retrieval.
 By evaluating performance differences between these two settings, we aim to understand how memory quality affects agent performance across the memory utilization stage for both tasks (single-memory tasks, joint-memory tasks).
 Performance degradation with lower-quality trajectories.
-Figure 11 shows the performance comparison between gold memory and retrieved memory across the memory utilization stage.
-In the memory utilization stage, high-capacity models (Llama-3.1-70b) show relatively stable performance across gold and retrieved memory, while lower-capacity models (Llama-3.1-8b) exhibit a substantial drop when using retrieved memory.
+Table 16 shows the performance comparison between high quality memory and acquired memory across the memory utilization stage.
+In the memory utilization stage, high-capacity models show relatively stable performance across gold and retrieved memory, while lower-capacity models exhibit a substantial drop when using retrieved memory.
 This suggests that less capable models have more difficulty extracting relevant information from an imperfect memory context.
-When executing more demanding joint-memory tasks, which require combining and reasoning across multiple memory sources, performance degrades sharply across all models with retrieved memory.
-This indicates that the challenge of integrating multiple memories introduces compounding complexity, as the agent must not only retrieve relevant information but also correctly synthesize disparate memory contexts.
-These findings emphasize a fundamental limitation of retrieval-based memory systems; while gold memory provides ground truth references, retrieved memory is inherently prone to semantic noise, as it relies on approximate similarity rather than precise matching.
-Enhancing memory quality through more precise filtering is essential to improve reasoning performance, particularly for complex joint tasks that require integration of multiple memory sources, and especially for smaller models with limited reasoning capabilities.
+When executing more demanding joint-memory tasks, which require combining and reasoning across multiple memory sources, performance degrades sharply across all models with acquired memory.
+This indicates that the compounded complexity of integrating multiple memories makes it increasingly important for the agent to rely on clear, noise-free, high-quality memory contexts from which synthesis can be performed reliably.
+Additionally, as shown in the table X, the overall reduction in planning steps when using high-quality memory suggests that episodic memory plays an important role in the planning process of embodied agents.
+Enhancing memory quality through more precise filtering is therefore essential not only for improving reasoning performance but also for increasing the agent’s planning efficiency.
 E.2.2 Robustness to Noisy and Partial Memories
@@
 This analysis provides insights into the relationship between retrieval precision and practical task execution in personalized embodied agents.
-Table 10: We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019).
+Table 18: We use an embedding-based retrieval method with the all-mpnet-base-v2 Sentence Transformer (Reimers and Gurevych, 2019).
 Retriever Strategy Top-K Missing Rates (%)
@@
 To assess whether the corresponding episodic memory is successfully retrieved during memory utilization, we first measure the missing rate under a baseline similarity-based retriever.
-[^1] We adopt an embedding-based method using the all-mpnet-base-v2 Sentence Transformer (Reimers & Gurevych, 2019).
+[^1] We adopt an embedding-based method using the all-mpnet-base-v2 Sentence Transformer (Reimers and Gurevych, 2019).
 The results show missing rates of 13.9% for top-3 and 3.5% for top-5, indicating that expanding to top-5 nearly eliminates missing cases.
@@
 Future work could build on these initial observations to better understand how agents transfer personalized knowledge across different contexts.
+E.4.4 User Profile Memory Robustness
+Experiment setup.
+In real-world interactions, users often refer to knowledge, objects, or locations using different names or expressions rather than clean and consistent symbolic identifiers.
+To directly evaluate the robustness of our proposed User Profile Memory under realistic conditions, we additionally conduct a qualitative analysis on how well User Profile Memory handles natural noise that arises in real user interactions.
+We identified two noise types that naturally arise when users interact with a personalized embodied agent:
+  • Knowledge Update Noise - the user expresses an existing knowledge using a different phrase (synonyms, paraphrases, alias names).
+  • Node Reference Noise - new knowledge refers to an existing object / location using a different form.
+For each category, three undergraduate annotators modified the original instructions grounded in the scene of Memento, producing 10 noisy instructions for each noise type.
+This procedure ensures the noise realistically reflects how humans describe object semantics or user patterns.
+These noisy instructions were crafted using the underlying scene (103997895_171031182103997895\_171031182), and to ensure an accurate evaluation of noise robustness, all experiments were conducted on graphs constructed from the original instructions.
+Full datasets are provided in Table 24 and Table 25
+Table 24: Dataset of Knowledge Update Noise Variants.
+Original Knowledge Modified Knowledge
+Gift teapot from grandfather Grandfather’s gifted teakettle
+Family heirloom vase Inherited family vase
+Gift toy airplane from childhood friend A toy airplane that a childhood friend gave me
+Favorite relaxation candle The candle I always use to relax
+Collectible android figure Android collectible figurine
+Bedroom organization Organizing setup for the bedroom space
+Morning work session setup Morning workspace arrangement
+Cozy room setup Arrangement for creating a cozy room atmosphere
+Bedtime setup Nighttime setup before going to bed
+Study materials organization Organized placement of study materials
+Table 25: Dataset of Node Reference Noise Variants.
+Original Description Modified Description
+Black teapot with curved spout and ornate lid A curved-spout black teapot featuring an intricately decorated lid
+Gray laptop with black keyboard/touchpad A gray laptop equipped with a black keyboard and touchpad
+Dark green analog alarm clock with twin bells Dark green twin-bell analog alarm clock
+Black and orange keychain with circular pendant A circular pendant keychain in black and orange
+android figure which is a black and white android panda figure a black and white Android panda figurine
+Move the toy airplane from the bedroom to the living room table. Place it next to the couch. Relocate the toy plane from the bedroom onto the table in the living room, positioning it next to the sofa.
+Move the picture frame and clock to the table in the bedroom.
+Then, place them next to each other on the counter in the kitchen.
+Bring the photo frame and the alarm clock to the bedroom table, and afterward arrange both items side by side on the countertop at kitchen.
+Place the bread, the potato, and the egg on the kitchen table and place them next to each other.
+Set the loaf, spud, and fresh egg together on the table at kitchen, arranging them close to one another.
+Move the laptop and the mouse pad to a new living room table.
+Move the clock to another shelf.
+Transfer the notebook computer and mouse mat onto the new table in the living room, and place the clock on a separate shelf.
+Take the book, pencil case, and multiport hub from the table in the living room and put them on the shelves in the bedroom, placing them next to each other.
+Gather the note, pencil pouch, and usb hub from the living room table and arrange them side by side on the bedroom shelves.
+Result.
+Table 26: Accuracy under different noise types.
+Noise Type Accuracy (%)
+Knowledge Update Noise 100.0
+Node Reference Noise 75.0
+[…]
+Table 27: Failure cases in node reference noise
+Existing Node Add Node
+mouse pad mouse mat
+clock alarm clock
+android figure which is a black and white android panda figure black and white Android Panda Figurine
+kitchen counter countertop at kitchen
+another shelf separate shelf
 Appendix F Prompts
@@
 F.3 Prompts for Memory Design Experiment
-Table 16: Episode metadata for the example task.
-Field Value
-episode_type user_pattern
-subtype preference
-task Could you decorate the bedroom as I like?
-task_percent_complete 1.0
-task_state_success 1.0
-Table 17: Trajectory for the example task.
-# Thought Action Result / State $\Delta$
-1 Since there are no objects found yet, I should start by exploring the bedroom to locate any objects that can be used for decoration. Explore
-(bedroom_1)
-No objects found
-2 Since no objects were found in bedroom_1, I should explore another room to locate objects that can be used for decoration.
-I will explore the living room next, as it is likely to contain decorative items.
-Navigate
-(to living_room_1)
-Location changed
-3 I have found a plant container and a statue on the shelves in the living room, which can be used for decoration in the bedroom. I will start by picking up the plant container from the shelves. Pick
-(plant_container_1)
-plant_container_1 held
-4 Now that I am close to the shelves, I will attempt to pick up the plant container to move it to the bedroom for decoration. (continue) Success
-5 I have successfully picked up the plant container. Now, I need to navigate to the bedroom to place it for decoration. Navigate
-(to bedroom_1)
-Arrived in bedroom_1
-6 I have successfully navigated to bedroom_1. Now, I need to find a suitable location to place the plant container for decoration. I will place it on the chest of drawers in the bedroom. Place
-(plant_container_1
-on chest_of_drawers_54)
-Object placed
-7 Now that I am close to chest_of_drawers_54, I will place the plant container on it for decoration. (continue) Confirmed placement
-8 I have successfully placed the plant container on chest_of_drawers_54 in bedroom_1.
-Now, I need to return to the living room to pick up the statue from the shelves for further decoration in the bedroom.
-Navigate
-(to living_room_1)
-Location changed
-9 I am back in the living room. I need to navigate closer to the shelves to pick up the statue for further decoration in the bedroom. Positioning for pick
-10 Now that I am close to the shelves, I will attempt to pick up the statue to move it to the bedroom for decoration. Pick
-(statue_0)
-statue_0 held
-11 I have successfully picked up the statue. Now, I need to navigate to the bedroom to place it for decoration. Navigate
-(to bedroom_1)
-Location changed
-12 I have successfully navigated to bedroom_1.
-Now, I need to find a suitable location to place the statue for decoration.
-I will place it on the chest of drawers next to the plant container.
-Positioning
-13 Now that I am close to chest_of_drawers_54, I will place the statue next to the plant container for decoration. Place
-(statue_0 on
-chest_of_drawers_54
-next_to plant_container_1)
-Object placed
-Table 18: Evaluation functions for the example task.
-# Function Args Description
-1 is_on_top object_handles=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], receptacle_handles=[bedroom_chest_of_drawers_handles],
-number=1
-Plant container is on drawer
-2 is_on_top object_handles=[Dog_Statue_4_:0000],
-receptacle_handles=[bedroom_chest_of_drawers_handles],
-number=1
-Statue is on drawer
-3 is_next_to entity_handles_a=[Cole_Hardware_Electric_Pot_Assortment_55_:0000], entity_handles_b=[Dog_Statue_4_:0000],
-number=1,
-l2_threshold=1.0
-Plant and statue are next to each other
-Table 19: Episode metadata for the example task.
-Field Value
-episode_type user_pattern
-subtype routine
-task Could you set up the candle for a cozy evening?
-task_percent_complete 0
-task_state_success 0
-Table 20: Trajectory for the example task.
-# Thought Action Result / State $\Delta$
-1 Since there are no objects found yet, I should start by exploring the tv room to locate a candle that can be set up for a cozy evening. Navigate
-(to tv_1)
-Location changed
-2 I have found a candle, candle_0, on table_34 in the tv room. To set it up for a cozy evening, I will first need to pick it up from the table. Pick
-(candle_0)
-candle_0 held
-3 Now that I am close to table_34, I will attempt to pick up the candle, candle_0, to set it up for a cozy evening. (continue) Success
-4 I have successfully picked up the candle, candle_0. To set it up for a cozy evening, I will place it on a suitable surface in the tv room, such as the tv table, table_14. Place
-(candle_0 on table_14)
-Object placed
-5 I am now close to table_14 in the tv room. I will place the candle, candle_0, on this table to set it up for a cozy evening. (continue) Failed to meet task constraints
-Table 21: Evaluation functions for the example task.
-# Function Args Description
-1 is_on_top object_handles=[candle], receptacle_handles=[chair_handles],
-number=1
-Candle must be on chair
-2 is_on_top object_handles=[candle], receptacle_handles=[table_handles],
-number=1
-Candle must be on table (after chair)
-3 Constraint TemporalConstraint: Placing the candle on the chair before placing it on the table.
-Figure 12: Example of a user profile memory graph representing object semantics knowledge.
-Figure 13: Example of a user profile memory graph representing user pattern knowledge.